### v1

# Proximality and selflessness for group C*-algebras

Abstract.
We prove that the reduced group $\mathrm{C}^{*}$-algebra of an infinite discrete group having a topologically-free extreme boundary is selfless in the sense of L. Robert.
Key words and phrases:
Extremely proximal actions, extreme boundaries, selfless $\mathrm{C}^{*}$-algebras
1991 Mathematics Subject Classification:
Primary 37B05; Secondary 46L35
The author was partially supported by JSPS KAKENHI Grant Numbers 24K00527, 25H00588, 25H00593
1. Introduction
A remarkable property called selflessness is recently introduced for a $\mathrm{C}^{*}$-algebra by L. Robert ([Ro]) and it quickly attracted a number of researchers’ attention (see, e.g., [AGKEP], [HKER], [RTV], [Vi]) as it implies many important regularity properties such as simplicity, (in the tracial setting) stable rank one, and strict comparison. Let’s say a group $\Gamma$ is $\mathrm{C}^{*}$-selfless if its reduced group $\mathrm{C}^{*}$-algebra is selfless. In this paper, we prove that every infinite countable discrete group with a minimal topologically-free extremely-proximal action is $\mathrm{C}^{*}$-selfless. Note that every $\mathrm{C}^{*}$-selfless group is $\mathrm{C}^{*}$-simple ([Ro]) and whether the converse is also true is an open problem (Problem XC in [STW]). Our result should be compared with Kalantar and Kennedy’s theorem ([KK]) that a discrete group $\Gamma$ is $\mathrm{C}^{*}$-simple if and only if it has a minimal topological-free strongly-proximal action. As the names suggest, strong proximality is weaker than extreme proximality ([Gl]).
Definition.
Let $\Gamma$ be a countable discrete group and $\Gamma\curvearrowright X$ be an action on a compact topological space $X$. The action $\Gamma\curvearrowright X$ (or the $\Gamma$-space $X$) is called an extreme boundary if it is minimal and extremely proximal ([Gl]) in the sense that for every non-empty open subsets $U$ and $V$ of $X$, there is $g\in\Gamma$ such that $g(X\setminus U)\subset V$.
For $x\in X$, we denote by $\Gamma_{x}\coloneq\{g\in\Gamma:gx=x\}$ the stabilizer subgroup at $x$. The action $\Gamma\curvearrowright X$ is said to be topologically-free if the set of points with trivial stabilizer group is dense in $X$. Note that if an extreme boundary $\Gamma\curvearrowright X$ is not topologically-free, then every proper closed subset of $X$ is pointwise fixed by some $g\in\Gamma\setminus\{1\}$.
We say a sequence $(z_{n})_{n}\in\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions. (1): For every neighborhoods $U_{\pm}$ of $z_{\pm}$ one has $z_{n}(X\setminus U_{-})\subset U_{+}$ eventually, or equivalently that $z_{n}^{-1}(X\setminus U_{+})\subset U_{-}$ eventually. (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$. Note that every topologically-free extreme boundary $X$ with $|X|>2$ accommodates an axial sequence.
Theorem 1.
An infinite countable discrete group $\Gamma$ having a topologically-free extreme boundary is $\mathrm{C}^{*}$-selfless. More precisely, for any axial sequence $(z_{n})_{n}$ in $\Gamma$ and any free ultrafilter $\mathcal{U}$, the homomorphism
$\Gamma*\mathopen{\langle}z\mathclose{\rangle}\to\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}},$
given by the constant embedding of $\Gamma$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ and $z\mapsto[\lambda(z_{n})]_{n}\in\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$, induces a faithful embedding of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(G*\mathopen{\langle}z\mathclose{\rangle})$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$.
See, e.g., [Bo], [BIO], [FLMMS], [IO], [JR], [LS], [LBMB] for examples of extreme boundaries. Every group satisfying the conclusion of Theorem 1 is mixed-identity-free ([HO]) and converse holds true if $\Gamma$ has a faithful minimal action of general type on a tree ([FLMMS]), since such an action gives rise to an extreme boundary. In particular, non-degenerate free product groups are $\mathrm{C}^{*}$-selfless. For the case of amalgamated free products and HNN extensions, see [FLMMS] and [IO] for the characterization of topological freeness of the corresponding Bass–Serre tree boundaries. Non-elementarily relatively hyperbolic groups ([Bo]) with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless, as their Bowditch compactifications are topologically-free extreme boundaries. It is not clear whether this generalizes to acylindrically hyperbolic groups, because their hyperbolic boundaries need not be compact.
Notations
A sequence is written as $(x_{n})_{n}$, while the equivalence class it represents in the ultrapower is written by $x_{n}]_{n}$. We use the symbol $z^{\pm}$ for a slightly abusive notation for “$z$ and/or $z^{-1}$” and similarly for $z_{\pm}$. For $\epsilon\in\{\pm\}$, we denote by $\bar{\epsilon}$ the opposite of $\epsilon$.
Acknowledgments
This research was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 4–8, 2025. He acknowledges the kind hospitality and the exciting environment provided by the institute. He thanks G. Patchell for the stimulating talk there and a useful comment on this paper.
2. Trees and covariant representations
We recall the notion of trees and their compactification. A tree is a connected graph without nontrivial cycles. See Section 5.2 and Appendix E in [BO] for more on this theme. Here we consider a countable tree $\mathrm{T}$ of infinite degree. The tree $\mathrm{T}$ is identified with its vertex set. A geodesic ray is an infinite sequence $\omega(0),\omega(1),\ldots$ in $\mathrm{T}$ such that $\omega(n)$ adjacent to $\omega(n-1)$ and $\omega(n)\neq\omega(n-2)$. Geodesic rays $\omega$ and $\omega^{\prime}$ are equivalent if there are $N$ and $N^{\prime}$ such that $\omega(N+n)=\omega^{\prime}(N^{\prime}+n)$ for all $n\geq 0$. The boundary $\partial\mathrm{T}$ is the equivalence classes of geodesic rays and $\bar{\mathrm{T}}\coloneq\mathrm{T}\cup\partial\mathrm{T}$ is the compactification of $\mathrm{T}$ (whose topology will be introduced shortly). We define $f\colon\mathrm{T}\times\bar{\mathrm{T}}\to\mathrm{T}$ as follows. For $s\in\mathrm{T}$, we set $f(s,s)=s$. For distinct $s,t\in\mathrm{T}$, we define $f(t,s)\in\mathrm{T}$ to be the unique point between $t$ and $s$ that is adjacent to $t$. For $t\in\mathrm{T}$ and $\omega\in\partial\mathrm{T}$, we define $f(t,\omega)\coloneq f(t,\omega(n))$ for any $n$ large enough. For each $t\in\mathrm{T}$, we consider a copy $\mathrm{T}_{t}$ of $\mathrm{T}$ and equip $\mathrm{T}_{t}$ a compact topology by taking the one-point compactification $\mathrm{T}\cup\{\infty\}$ of $\mathrm{T}$ and identify the new point $\infty$ with $t$. That is, every sequence in $\mathrm{T}$ that diverges to infinity converges to $t$ in $\mathrm{T}_{t}$. We consider the embedding
$\bar{\mathrm{T}}\ni\omega\mapsto(f(t,\omega))_{t}\in\prod_{t\in\mathrm{T}}\mathrm{T}_{t}$
and equip $\bar{\mathrm{T}}$ with the induced topology. This makes $\bar{\mathrm{T}}$ a second-countable compact topological space (Proposition 5.2.5 in [BO]). Note that if $s\in\mathrm{T}$ and $t_{n}\in\mathrm{T}$ are such that $f(s,t_{n})\to\infty$ in $\mathrm{T}$, then $t_{n}\to s$ in $\bar{\mathrm{T}}$.
Theorem 2.
Let $\Lambda$ be a countable discrete group acting on a countable tree $\mathrm{T}$ and on a unital $\mathrm{C}^{*}$-algebra $A$. Let $\pi\colon\Lambda\ltimes_{\max}A\to\mathbb{B}(\mathcal{H})$ be a covariant representation. Assume that there is a $\Lambda$-equivariant unital positive map from $C(\bar{\mathrm{T}})$ into the center of $A$ and that for every $t]\in\mathrm{T}/\Lambda$, the restriction of $\pi$ to the stabilizer subgroup $\Lambda_{t}\coloneq\{g\in\Lambda:gt=t\}$ is continuous on the reduced crossed product $\Lambda_{t}\ltimes_{\min}A$. Then $\pi$ is continuous on the reduced crossed product $\Lambda\ltimes_{\min}A$
Proof.
Let $\pi\colon\Lambda\ltimes_{\max}A\to\mathbb{B}(\mathcal{H})$ be given. We denote by $\lambda$ the unitary representation of $\Lambda$ on $\ell_{2}\mathrm{T}$ induced by $\Lambda\curvearrowright\mathrm{T}$ and denote by $\sigma$ the $\Lambda$ action on $A$. We claim that $\pi$ is weakly contained in the representation $\tilde{\pi}$ on $\ell_{2}\mathrm{T}\otimes\mathcal{H}$ given by $g\mapsto(\lambda\otimes\pi)(g)$ for $g\in\Lambda$ and $a\mapsto 1\otimes\pi(a)$ for $a\in A$. Indeed, there is a sequence $\zeta_{n}\colon\bar{\mathrm{T}}\to\mathrm{Prob}(\mathrm{T})$ of Borel maps such that $\sup_{x\in\bar{\mathrm{T}}}\|g\zeta_{n}^{x}-\zeta_{n}^{gx}\|_{1}\to 0$ for every $g\in\Lambda$ (Lemma 5.2.6 in [BO]). It follows that $h_{n}^{t}(x)\coloneq\zeta_{n}^{x}(t)$ defines sequences of positive Borel functions on $\bar{\mathrm{T}}$ such that $\sum_{t\in\mathrm{T}}h_{n}^{t}=1$ and
$\|\sum_{t\in\mathrm{T}}|h_{n}^{gt}-\sigma_{g}(h_{n}^{t})|\|_{\infty}\to 0.$
NB: the above sum is an infinite sum of positive Borel functions. Let $\theta$ denote the $\Lambda$-equivariant unital completely positive map from $C(\bar{\mathrm{T}})$ into the center of $A$. The map $\pi\circ\theta$ canonically extends to a normal unital positive map $\Theta$ from the $\mathrm{C}^{*}$-algebra $B(\bar{\mathrm{T}})$ of the bounded Borel functions into $(\pi\circ\theta)(C(\bar{\mathrm{T}}))^{\prime\prime}$. By “normal”, we mean that it sends a bounded pointwise convergent sequence to a strong operator topology convergent sequence. We note that $\Theta$ satisfies $|\Theta(f)|\leq\Theta(|f|)$ because the range of $\Theta$ is commutative. We define isometries $V_{n}\colon\mathcal{H}\to\ell_{2}\mathrm{T}\otimes\mathcal{H}$ by $V_{n}\xi\coloneq\sum_{t\in\mathrm{T}}\delta_{t}\otimes\Theta(h_{n}^{t})^{1/2}\xi$, or $V_{n}$ is the column operator $(\Theta(h_{n}^{t})^{1/2})_{t\in\mathrm{T}}$. Then
$\|V_{n}\pi(g)-(\lambda\otimes\pi)(g)V_{n}\|^{2}=\|\sum_{t}|\Theta(h_{n}^{gt})^{1/2}-\Theta(\sigma_{g}(h_{n}^{t}))^{1/2}|^{2}\|_{\infty}\to 0$
for every $g\in\Lambda$ and $V_{n}\pi(a)=(1\otimes\pi(a))V_{n}$ for $a\in A$. This proves the claim that $\pi$ is weakly contained in $\tilde{\pi}$. Thus continuity of $\pi$ follows from that of $\tilde{\pi}$.
For the continuity of $\tilde{\pi}$, it suffices to consider the continuity of the vector states $\varphi$ associated with the vectors of the form $\delta_{t}\otimes\xi$, since they form a cyclic family. But $\varphi=\psi\circ E$, where $E$ is the conditional expectation from $\Lambda\ltimes_{\max}A$ onto $\Lambda_{t}\ltimes_{\max}A$ and $\psi$ is the vector state associated with $\xi$. (We do not need the fact that the natural embedding $\Lambda_{t}\ltimes_{\max}A\subset\Lambda\ltimes_{\max}A$ is faithful.) Hence the continuity on the reduced crossed product $\Lambda\ltimes_{\min}A$ follows from that for $\Lambda_{t}\ltimes_{\min}A$ for $t\in\mathrm{T}$ one from each $\Lambda$-orbit. ∎
3. A space with “tree structure”
Let $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. By the normal form of $g\in\tilde{\Gamma}$, we mean the expression
$g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots z^{\epsilon_{l}}a_{l+1}$
with $a_{i}\in\Gamma$ and $\epsilon_{i}\in\{\pm\}$, which is reduced in the sense that if $i>1$ and $a_{i}=1$, then $\epsilon_{i-1}=\epsilon_{i}$.
We introduce the tree $\mathrm{T}\coloneq\tilde{\Gamma}/\Gamma$ on which $\tilde{\Gamma}$ acts. We declare that the points $s,t\in\mathrm{T}$ are adjacent if $s^{-1}t\in\{\Gamma z^{\pm}\Gamma\}$. With this graph structure $\mathrm{T}$ becomes a tree (of infinite degree). We call the point $o\coloneq\Gamma\in\mathrm{T}$ the origin. Every $t\in\mathrm{T}$ has unique representative $\rho(t)\in\tilde{\Gamma}$ whose normal form ends by $z^{\pm}$, or $\rho(o)=1$ if $t=o$. Note that $\rho(at)=a\rho(t)$ for all $a\in\Gamma$ and $t\in\mathrm{T}\setminus\{o\}$ and that $\rho(z^{\pm}t)=z^{\pm}\rho(t)$ for all $t\in\mathrm{T}$.
Suppose first that $s$ and $t$ are adjacent. We label the oriented edge from $t$ to $s$ by $\ell(t,s)\coloneq\kappa(\rho(t)^{-1}s)\in\Gamma\{z_{\pm}\}$, where $\kappa(az^{\pm}\Gamma)\coloneq az_{\pm}$. Note that $t$ and $\ell(t,s)$ uniquely determines $s$. E.g., $\ell(az\Gamma,azbz\Gamma)=bz_{+}$ and $\ell(azbz\Gamma,az\Gamma)=z_{-}$ for every $a,b\in\Gamma$. The Figure 1 shows how $\mathrm{T}$ and $\ell$ look like. Next we extend the label on $(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$ by setting $\ell(t,s)\coloneq\ell(t,t_{1})$, where $t_{1}$ is the unique point between $t$ and $s$ that is adjacent to $t$, or colloquially, $\ell(t,s)$ is the first $\Gamma\{z^{\pm}\}$ component in the normal form $\rho(t)^{-1}\rho(s)$. For example, $\ell(az,azbzcz\Gamma)=bz_{+}$ and $\ell(cz\Gamma,azbz\Gamma)=z_{-}$ unless $a=c$.
Figure 1. The tree $\mathrm{T}$
The following is not hard to see.
Lemma 3.
Let $(t,s)\in(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$. For $a\in\Gamma$, one has $\ell(at,as)=\ell(t,s)$ for $t\neq o$ and $\ell(o,as)=a\ell(o,s)$. Also, $\ell(zt,zs)=\ell(t,s)$.
Let $\Gamma\curvearrowright X$ be a topologically-free extreme boundary with an axial sequence $(z_{n})_{n}$. We denote by $\varphi_{n}\colon\tilde{\Gamma}\to\Gamma$ the homomorphism that is identity on $\Gamma$ and send $z$ to $z_{n}\in\Gamma$.
Lemma 4.
Let $g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots z^{\epsilon_{l}}$ be an element in the normal form with $l\geq 1$ and $x\in X\setminus\{z_{\bar{\epsilon}_{l}}\}$. Then $\varphi_{n}(g)x\to a_{1}z_{\epsilon_{1}}$ as $n\to\infty$.
Proof.
Let $U_{\pm}$ be arbitrarily small neighborhoods of $z_{\pm}$ that satisfy $x\notin U_{\bar{\epsilon}_{l}}$ and $a_{i}(U_{+}\cup U_{-})\cap(U_{+}\cup U_{-})=\emptyset$. Then by the ping-pong argument, one has $\varphi_{n}(g)x\in a_{1}U_{\epsilon_{1}}$ for all $n$ large enough. ∎
We denote by $\sigma$ the corresponding action of $\Gamma$ on $C(X)\subset C(X)^{\mathcal{U}}$, where the embedding is the constant embedding. We denote by $\sigma_{z}$ the automorphism arising from the sequence $(\sigma_{z_{n}})_{n}$. They together give rise to the action of $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$ on $C(X)^{\mathcal{U}}$, which is given by
$\sigma_{g}([f_{n}]_{n})=[\sigma_{\varphi_{n}(g)}(f_{n})]_{n}.$
We are interested in the commutative $\mathrm{C}^{*}$-algebra
$C(X_{\mathrm{T}})\coloneq\mathrm{C}^{*}(\bigcup_{g\in\tilde{\Gamma}}\sigma_{g}(C(X)))\subset C(X)^{\mathcal{U}}$
on which $\tilde{\Gamma}$ acts, where $X_{\mathrm{T}}$ is the Gelfand spectrum of $C(X_{\mathrm{T}})$. By definition, there is a surjective homomorphism
$Q\colon C(X^{\mathrm{T}})\cong\bigotimes_{t\in\mathrm{T}}C(X)\to C(X_{\mathrm{T}}),\quad Q(f^{(s)})=\sigma_{\rho(s)}(f),$
where $f^{(s)}\in C(X^{\mathrm{T}})$ is defined for $f\in C(X)$ and $\mathbf{x}=(x_{t})_{t}$ by $f^{(s)}(\mathbf{x})\coloneq f(x_{s})$. This gives rise to an embedding $Q_{*}\colon X_{T}\to X^{\mathrm{T}}$. We equip $X^{\mathrm{T}}$ a $\tilde{\Gamma}-$action by declaring that $a\in\Gamma$ acts at $\mathbf{x}=(x_{t})_{t}\in X^{\mathrm{T}}$ by $(a\mathbf{x})_{o}\coloneq ax_{o}$ and $(a\mathbf{x})_{t}\coloneq x_{a^{-1}t}$ for $t\in\mathrm{T}\setminus\{o\}$, and that $z$ acts by $(z\mathbf{x})_{t}\coloneq x_{z^{-1}t}$.
Lemma 5.
The embedding $Q_{*}\colon X_{\mathrm{T}}\to X^{\mathrm{T}}$ is $\tilde{\Gamma}-$equivariant.
Proof.
Let $\varsigma$ denote the corresponding $\tilde{\Gamma}-$action on $C(X^{\mathrm{T}})$. It is enough to check that
$Q(\varsigma_{a}(f^{(o)}))=Q(\sigma_{a}(f)^{(o)})=\sigma_{a}(f)=\sigma_{a}(Q(f^{(o)})),$
$Q(\varsigma_{a}(f^{(s)}))=Q(f^{(as)})=\sigma_{\rho(as)}(f)=\sigma_{a}(\sigma_{\rho(s)}(f))=\sigma_{a}(Q(f^{(s)}))$
for $s\neq o$ and,
$Q(\varsigma_{z}(f^{(s)}))=Q(f^{(zs)})=\sigma_{\rho(zs)}(f)=\sigma_{z}(\sigma_{\rho(s)}(f))=\sigma_{z}(Q(f^{(s)})).$
Hereafter, we omit writing $Q$ and identify $\varsigma$ with $\sigma$. Thus we view $X_{\mathrm{T}}$ as a compact subset of $X^{\mathrm{T}}$. For every $s\in\mathrm{T}$, we define the continuous embedding $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ by $\theta(s,x)_{s}\coloneq x$ and $\theta(s,x)_{t}\coloneq\ell(t,s)$ for $t\neq s$. We also define the continuous embedding $\theta\colon\partial\mathrm{T}\to X^{\mathrm{T}}$ by $\theta(\omega)_{t}\coloneq\ell(t,\omega)$. Note that $\omega\to\theta(\omega)_{t}$ is locally constant.
The following theorem asserts that $\Gamma\curvearrowright X_{\mathrm{T}}$ has a “tree structure”. See Figure 2.
Figure 2. The space $X_{\mathrm{T}}$ around $\theta(s,X)$, where $s=az\Gamma$ and $t=azbz\Gamma$, $\theta(s,bz_{+})=\theta(t,z_{-})$ and $\theta(s,z_{-})=\theta(o,az)$.
Theorem 6.
In the above setting for $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$, the following hold.
  1. (1)
    $X_{\mathrm{T}}=\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})$.
  2. (2)
    $\theta(\mathrm{T},X)\cap\theta(\partial\mathrm{T})=\emptyset$.
  3. (3)
    For $s\neq t$ and $x,y\in X$, one has $\theta(s,x)=\theta(t,y)$ if and only if $s$ and $t$ are adjacent and $x=\ell(s,t)$ and $y=\ell(t,s)$.
  4. (4)
    For $a\in\Gamma$, one has $a\theta(o,x)=\theta(o,ax)$ and $a\theta(s,x)=\theta(as,x)$ for $s\neq o$.
  5. (5)
    $z\theta(s,x)=\theta(zs,x)$.
  6. (6)
    $\theta\colon\partial\Gamma\to X_{\mathrm{T}}$ is $\tilde{\Gamma}-$equivariant.
  7. (7)
    A sequence $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(s,x)$ in $X_{\mathrm{T}}$ if and only if either $(t_{n})_{n}$ converges to $s$ in $\bar{\mathrm{T}}$ and the sequence $(y_{n})_{n}$, defined by $y_{n}=x_{n}$ if $t_{n}=s$ and $y_{n}=\ell(s,t_{n})$ if $t_{n}\neq s$, converges to $x$ in $X$; or there is an adjacent point $s_{1}$ to $s_{0}\coloneq s$ in $\mathrm{T}$ such that $t_{n}\in\{s_{0},s_{1}\}$ eventually, $x=\ell(s_{0},s_{1})$, and the subsequences $(x_{n})_{\{n:t_{n}=s_{i}\}}$ (possibly there is only one of them) converge to $\ell(s_{i},s_{j})$ for $\{i,j\}=\{0,1\}$. A sequence $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(\omega)$ in $X_{\mathrm{T}}$ if and only if $(t_{n})_{n}$ converges to $\omega$ in $\bar{\mathrm{T}}$.
  8. (8)
    A sequence $\theta(\omega_{n})$ converges to $\theta(s,x)$ in $X_{\mathrm{T}}$ if and only if $(\omega_{n})_{n}$ converges to $s$ in $\bar{\mathrm{T}}$ and and $(\ell(s,\omega_{n}))_{n}$ converges to $x$ in $X$.
  9. (9)
    One has $\theta(\mathrm{T},\Gamma\{z_{\pm}\})=\tilde{\Gamma}\theta(o,\{z_{\pm}\})$. The map $\beta\colon X_{\mathrm{T}}\setminus\tilde{\Gamma}\theta(o,\{z_{\pm}\})\to\bar{\mathrm{T}}$, given by $\beta(\theta(s,x))=s$ and $\beta(\theta(\omega))=\omega$, is $\tilde{\Gamma}-$equivariant and continuous.
Proof.
[…]
The assertions (4-6) follow from Lemma 3; E.g., for $a\in\Gamma$ and $s\in\mathrm{T}\setminus\{o\}$, one has
$(a\theta(s,x))_{o}=a\theta(s,x)_{o}=a\ell(o,s)=\ell(o,as)=\theta(as,x)_{o}$
and for $t\neq o$
$(a\theta(s,x))_{t}=\theta(s,x)_{a^{-1}t}=\left\{\begin{array}[]{ll}x&\mbox{ if $t=as$}\\ \ell(a^{-1}t,s)=\ell(t,as)&\mbox{ if $t\neq as$}\end{array}\right\}=\theta(as,x)_{t}.$
The “if” part of (7-8) are easy. For the “only if” part, suppose that $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(s,x)$. If $(t_{n})_{n}$ is converges to some point $s^{\prime}$, then by the “if” part and the compactness of $X$, one has $\theta(t_{n},x_{n})\to\theta(s^{\prime},y)$, where $y$ is the limit of $y_{n}\coloneq x_{n}$ if $t_{n}=s^{\prime}$ and $y_{n}\coloneq\ell(s^{\prime},t_{n})$ if $t_{n}\neq s^{\prime}$. Since $\theta(s^{\prime},y)=\theta(s,x)$, one has either $s^{\prime}=s$ or $s^{\prime}$ is an adjacent point to $s$ that satisfies $x=\ell(s,s^{\prime})$ and $y=\ell(s^{\prime},s)$. This proves the “only if” part for the case $(t_{n})_{n}$ is convergent. Now consider the case where $(t_{n})_{n}$ is not convergent. By the previous case, the only limit points are either $s$ or the unique adjacent point $s^{\prime}$ that satisfies $x=\ell(s,s^{\prime})$. The other claimed conditions also follow. The rest of the proof of (7-8) is similar.
We prove (1). We first prove that $\theta(\mathrm{T},X)\subset X_{\mathrm{T}}$. Let $s\in\mathrm{T}$ and $x\in X\setminus\Gamma\{z_{\pm}\}$ be given. We consider the character $\chi$ on $C(X_{\mathrm{T}})\subset C(X)^{\mathcal{U}}$ arising from the sequence $(\varphi_{n}(\rho(s))x)_{n}$. Then by Lemma 4 one has
$\chi(f^{(t)})=\lim\nolimits_{\mathcal{U}}\sigma_{\varphi_{n}(\rho(t))}(f)(\varphi_{n}(\rho(s))x)=\lim\nolimits_{\mathcal{U}}f(\varphi_{n}(\rho(t)^{-1}\rho(s))x)=f(\ell(t,s))$
for $t\neq s$ and $\chi(f^{(s)})=f(x)$. This means $\chi$ corresponds to the point $\theta(s,x)$ in $X_{\mathrm{T}}$. Since $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ is continuous and $X\setminus\Gamma\{z_{\pm}\}$ is dense in $X$, one sees that $\theta(s,X)\subset X_{\mathrm{T}}$. Similarly for $\theta(\partial\mathrm{T})$. It follows that $\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})\subset X_{\mathrm{T}}$.
It remains to prove the opposite inclusion. Let $\mathbf{x}\coloneq(x_{t})_{t}\in X^{\mathrm{T}}$ be given and assume that $\mathbf{x}\in X_{\mathrm{T}}$. We define $w\colon\mathrm{T}\to\mathrm{T}$, “the orientation of $\mathbf{x}$ at tt”, as follows. If $x_{t}\in X\setminus\Gamma\{z_{\pm}\}$, then $w(t)\coloneq t$. If $x_{t}\in\Gamma\{z_{\pm}\}$, then $w(t)\coloneq\rho(t)\kappa^{-1}(x_{t})$, where $\kappa^{-1}(az_{\pm})\coloneq az^{\pm}\Gamma\in\mathrm{T}$. Note that $w(t)$ is adjacent to $t$ in the latter case. We claim that if $s,t\in\mathrm{T}$ are adjacent, then it cannot happen that $w(s)\neq t$ and $w(t)\neq s$ simultaneously. Suppose for a contradiction that it happened. We may assume that $s$ is between $o$ and $t$ and $\rho(t)=\rho(s)az$ with no cancellation. Thus one has $x_{s}\neq az_{+}$ and $x_{t}\neq z_{-}$. Let $f\in C(X)$ be such that $f(x_{s})=1$ and $f=0$ on a neighborhood of $az_{+}$. Similarly let $g$ be such that $g(x_{t})=1$ and $g=0$ on a neighborhood of $z_{-}$. We consider the function $f^{(s)}g^{(t)}$ on $X^{\mathrm{T}}$. It satisfies $(f^{(s)}g^{(t)})(\mathbf{x})=f(x_{s})g(x_{t})=1$. However,
$(f^{(s)}g^{(t)})|_{X_{\mathrm{T}}}=Q(f^{(s)}g^{(t)})=\sigma_{\rho(s)}(f)\sigma_{\rho(t)}(g)=\sigma_{\rho(s)}(f\sigma_{az}(g))=0$
since $(\operatorname{supp}f)\cap az_{n}(\operatorname{supp}g)=\emptyset$ eventually. This means that $\mathbf{x}\in X^{\mathrm{T}}\setminus X_{\mathrm{T}}$, a contradiction. Thus we have proved the claim for adjacent points $s,t\in\mathrm{T}$ that $w(s)\neq t$ implies $w(t)=s$. This yields three possibilities. (1): There is $s\in\mathrm{T}$ such that $w$ is oriented to the $s$ direction. (2): There is $\omega\in\partial\mathrm{T}$ such that $w$ is oriented to the $\omega$ direction. (3): There is a pair $s_{1},s_{2}$ of adjacent points such that $w$ is oriented to $\{s_{1},s_{2}\}$ direction and $w(s_{i})=s_{j}$ for $\{i,j\}=\{1,2\}$. In the first case, one has $\mathbf{x}=\theta(s,x_{s})$. In the second case, one has $\mathbf{x}=\theta(\omega)$. In the third case, one has $\mathbf{x}=\theta(s_{1},\ell(s_{1},s_{2}))=\theta(s_{2},\ell(s_{2},s_{1}))$. It follows that $\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})\supset X_{\mathrm{T}}$.
The assertion (9) is a consequence of (1-8). ∎
4. Proof of Theorem 1
Proof of Theorem 1.
By Theorem 2.6 in [Ro], $\mathrm{C}^{*}$-selflessness follows from the “more precisely” statement. Let $\Gamma\curvearrowright X$ be a topologically-free extreme boundary with an axial sequence $(z_{n})_{n}$ and $\pi\colon\Gamma*\mathopen{\langle}z\mathclose{\rangle}\to\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ be the corresponding homomorphism. We set $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. We take a $\Gamma$-equivariant embedding $C(X)\subset\ell_{\infty}\Gamma\subset\mathbb{B}(\ell_{2}\Gamma)$. It gives rise to $\tilde{\Gamma}-$equivariant embedding of $C(X_{\mathrm{T}})$ into $\mathbb{B}(\ell_{2}\Gamma)^{\mathcal{U}}$, where $C(X_{\mathrm{T}})$ is defined in the previous section. We take a faithful representation of $\mathbb{B}(\ell_{2}\Gamma)^{\mathcal{U}}$ on a Hilbert space $\mathcal{H}$ and view $\pi$ as a covariant representation of $\tilde{\Gamma}\ltimes C(X_{\mathrm{T}})$. Then the representation $\pi|_{C(X_{\mathrm{T}})}$ canonically extends to a normal representation $\pi$ of the $\mathrm{C}^{*}$-algebra $B(X_{\mathrm{T}})$ of the bounded Borel functions on $X_{\mathrm{T}}$. Let $p$ denote the orthogonal projection on $\mathcal{H}$ corresponding to the countable subset $\tilde{\Gamma}\{z_{\pm}\}$ of $X_{\mathrm{T}}$, which is possibly zero. The unitary representation $p\pi$ of $\tilde{\Gamma}$ is covariant with $\ell_{\infty}\tilde{\Gamma}$ and hence unitarily equivalent to a multiple of the regular representation. By Theorem 6, there is a $\tilde{\Gamma}-$equivariant embedding of $C(\bar{\mathrm{T}})$ into $p^{\perp}\pi(B(X_{\mathrm{T}}))$. Thus by Theorem 2 and the fact that $\pi|_{\Gamma}$ is weakly equivalent to the regular representation of $\Gamma$, the unitary representation $p^{\perp}\pi$ of $\tilde{\Gamma}$ is weakly contained in the regular representation. ∎
A crossed product analogue of Theorem 1 holds true under the additional assumption that $\Gamma$ is exact, which ensures $\Gamma\ltimes_{\min}A^{\mathcal{U}}\subset(\Gamma\ltimes_{\min}A)^{\mathcal{U}}$.

### v6

@@
 Abstract.
-We prove that the reduced group $\mathrm{C}^{*}$-algebra of an infinite discrete group having a topologically-free extreme boundary is selfless in the sense of L. Robert.
+We prove that the reduced group $\mathrm{C}^{*}$-algebras of infinite countable discrete groups having topologically-free extreme boundaries, or more generally groups that satisfy certain combinatorial property including all acylindrically hyperbolic groups with no nontrivial finite normal subgroups and all Zariski-dense subgroups of $\mathrm{PSL}(n,\mathbb{R})$, are selfless in the sense of L.
+Robert.
+This generalizes the recent results of Amrutam, Gao, Kunnawalkam Elayavalli, and Patchell, and of Vigdorovich.
+We also prove that selflessness is stable under tensor product among exact $\mathrm{C}^{*}$-algebras and that a $\mathrm{C}^{*}$-probability space is selfless provided that it is either simple and purely infinite or simple, exact, $\mathcal{Z}$-stable, and uniquely tracial.
 Key words and phrases:
@@
 Robert ([Ro]) and it quickly attracted a number of researchers’ attention (see, e.g., [AGKEP], [HKER], [RTV], [Vi]) as it implies many important regularity properties such as simplicity, (in the tracial setting) stable rank one, and strict comparison.
-Let’s say a group $\Gamma$ is $\mathrm{C}^{*}$-selfless if its reduced group $\mathrm{C}^{*}$-algebra is selfless.
+In particular, Amrutam, Gao, Kunnawalkam Elayavalli, and Patchell ([AGKEP]) has proved that a large family of groups, namely acylindrically hyperbolic groups with the rapid decay property and with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless.
+Here, we say a group $\Gamma$ is $\mathrm{C}^{*}$-selfless if it its reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(\Gamma)$ is selfless.
 In this paper, we prove that every infinite countable discrete group with a minimal topologically-free extremely-proximal action is $\mathrm{C}^{*}$-selfless.
-Note that every $\mathrm{C}^{*}$-selfless group is $\mathrm{C}^{*}$-simple ([Ro]) and whether the converse is also true is an open problem (Problem XC in [STW]).
-Our result should be compared with Kalantar and Kennedy’s theorem ([KK]) that a discrete group $\Gamma$ is $\mathrm{C}^{*}$-simple if and only if it has a minimal topological-free strongly-proximal action.
+Our method is topological, as opposed to the more analytical one in [AGKEP], and to construct a kind of “tree-graded” space ([DS]) out of an extreme boundary (see Section 3).
+Note that every $\mathrm{C}^{*}$-selfless group is $\mathrm{C}^{*}$-simple ([Ro]) and whether the converse is also true is an open problem (Problem XCI in [STW]).
+We remind Kalantar and Kennedy’s theorem ([KK]) that a discrete group is $\mathrm{C}^{*}$-simple if and only if it has a minimal topological-free strongly-proximal action.
 As the names suggest, strong proximality is weaker than extreme proximality ([Gl]).
@@
 For $x\in X$, we denote by $\Gamma_{x}\coloneq\{g\in\Gamma:gx=x\}$ the stabilizer subgroup at $x$.
-The action $\Gamma\curvearrowright X$ is said to be topologically-free if the set of points with trivial stabilizer group is dense in $X$.
+The action $\Gamma\curvearrowright X$ is said to be topologically-free if the set of points with trivial stabilizer groups is dense in $X$.
 Note that if an extreme boundary $\Gamma\curvearrowright X$ is not topologically-free, then every proper closed subset of $X$ is pointwise fixed by some $g\in\Gamma\setminus\{1\}$.
-We say a sequence $(z_{n})_{n}\in\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions.
+We say a sequence $(z_{n})_{n}$ in $\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions.
 (1): For every neighborhoods $U_{\pm}$ of $z_{\pm}$ one has $z_{n}(X\setminus U_{-})\subset U_{+}$ eventually, or equivalently that $z_{n}^{-1}(X\setminus U_{+})\subset U_{-}$ eventually.
 (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$.
-Note that every topologically-free extreme boundary $X$ with $|X|>2$ accommodates an axial sequence.
+Note that an extreme boundary $X$ with $|X|>2$ accommodates an axial sequence if and only if it is topologically-free.
 Theorem 1.
 An infinite countable discrete group $\Gamma$ having a topologically-free extreme boundary is $\mathrm{C}^{*}$-selfless.
-More precisely, for any axial sequence $(z_{n})_{n}$ in $\Gamma$ and any free ultrafilter $\mathcal{U}$, the homomorphism
+More precisely, for any axial sequence $(z_{n})_{n}$ in $\Gamma$ and any free ultrafilter $\mathcal{U}$ on $\mathbb{N}$, the homomorphism
 $\Gamma*\mathopen{\langle}z\mathclose{\rangle}\to\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}},$
-given by the constant embedding of $\Gamma$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ and $z\mapsto[\lambda(z_{n})]_{n}\in\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$, induces a faithful embedding of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(G*\mathopen{\langle}z\mathclose{\rangle})$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$.
+given by the diagonal embedding of $\Gamma$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ and $z\mapsto[z_{n}]_{n}\in\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$, induces a faithful embedding of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(G*\mathopen{\langle}z\mathclose{\rangle})$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$.
 See, e.g., [Bo], [BIO], [FLMMS], [IO], [JR], [LS], [LBMB] for examples of extreme boundaries.
-Every group satisfying the conclusion of Theorem 1 is mixed-identity-free ([HO]) and converse holds true if $\Gamma$ has a faithful minimal action of general type on a tree ([FLMMS]), since such an action gives rise to an extreme boundary.
-In particular, non-degenerate free product groups are $\mathrm{C}^{*}$-selfless.
-For the case of amalgamated free products and HNN extensions, see [FLMMS] and [IO] for the characterization of topological freeness of the corresponding Bass–Serre tree boundaries.
-Non-elementarily relatively hyperbolic groups ([Bo]) with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless, as their Bowditch compactifications are topologically-free extreme boundaries.
-It is not clear whether this generalizes to acylindrically hyperbolic groups, because their hyperbolic boundaries need not be compact.
+Every group satisfying the conclusion of Theorem 1 is mixed-identity-free (see [HO] for the terminology) and converse holds true if $\Gamma$ has a faithful minimal action of general type on a tree ([FLMMS]), since such an action gives rise to an extreme boundary.
+In particular, non-elementary free product groups are $\mathrm{C}^{*}$-selfless.
+For the case of amalgamated free products and HNN extensions, see [FLMMS] and [IO] for a characterization of topological freeness of the corresponding Bass–Serre tree compactifications.
+Non-elementarily relatively hyperbolic groups with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless, as their Bowditch compactifications ([Bo]) are topologically-free extreme boundaries.
+It is not clear whether this generalizes to acylindrically hyperbolic groups (with no assumption on the rapid decay property, cf.
+[AGKEP]), because their hyperbolic structure need not yield suitable compactifications.
+We fix this problem in Section 8 and prove that a group satisfying property $\mathrm{P}_{\mathrm{PHP}}$ (defined there) is $\mathrm{C}^{*}$-selfless.
+The class of groups with property $\mathrm{P}_{\mathrm{PHP}}$ contains all acylindrically hyperbolic groups with no nontrivial finite normal subgroups as well as all Zariski-dense subgroups of $\mathrm{PSL}(d\geq 2,\mathbb{R})$ (cf.
+[Vi]).
+There are $\mathrm{C}^{*}$-selfless groups outside the mixed-identity-free realm.
+Theorem 2.
+Let $(A_{i},\varphi_{i})$ be separable $\mathrm{C}^{*}$-probability spaces.
+Assume that all $(A_{i},\varphi_{i})$ are selfless and exact.
+Then the tensor product $\bigotimes_{i}(A_{i},\varphi_{i})$ is selfless.
+In particular, the class of countable discrete groups that are $\mathrm{C}^{*}$-selfless and exact is closed under direct product.
+We introduce the notion of complete selflessness in Section 6 and prove that the examples in Theorem 1 and their tensor product are in fact completely selfless, while no nuclear tracial $\mathrm{C}^{*}$-probability space is completely selfless.
+The following is a partial converse to Theorem 3.1 in [Ro] and solves Question 5.4 in [Ro].
+Note that by Matui and Sato’s theorem ([MS]), $\mathcal{Z}$-stability is necessary for a nuclear tracial $\mathrm{C}^{*}$-probability space to be selfless.
+Theorem 3.
+A simple and purely infinite $\mathrm{C}^{*}$-probability space $(A,\varphi)$ is completely selfless.
+A simple, exact, $\mathcal{Z}$-stable, and uniquely tracial $\mathrm{C}^{*}$-probability space $(A,\tau)$ is selfless.
 Notations
-A sequence is written as $(x_{n})_{n}$, while the equivalence class it represents in the ultrapower is written by $x_{n}]_{n}$.
-We use the symbol $z^{\pm}$ for a slightly abusive notation for “$z$ and/or $z^{-1}$” and similarly for $z_{\pm}$.
-For $\epsilon\in\{\pm\}$, we denote by $\bar{\epsilon}$ the opposite of $\epsilon$.
+We use the symbol $\pm$” for a slightly abusive notation for “++ and/or -” or “+1+1 and/or 1-1”.
+We denote by $A^{\mathcal{U}}$ the ultrapower of a $\mathrm{C}^{*}$-algebra $A$ and by $a_{n}]_{n}$ the element in $A^{\mathcal{U}}$ represented by a sequence $(a_{n})_{n}$ in $A$.
+A $\mathrm{C}^{*}$-probability space is a pair $(A,\varphi)$ of a unital $\mathrm{C}^{*}$-algebra $A$ and a distinguished state $\varphi$ which is not necessarily faithful but assumed GNS-faithful (a.k.a.
+non-degenerate) unless the $\mathrm{C}^{*}$-algebra in consideration is an ultrapower or an ultraproduct.
 Acknowledgments
-This research was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 4–8, 2025.
-He acknowledges the kind hospitality and the exciting environment provided by the institute.
+A part of this work was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 3–8, 2025 and the Isaac Newton Institute for the program “Operators, Graphs, Groups” in October 2025.
+He acknowledges the kind hospitality and the exciting environment provided by the institutes.
 He thanks G.
-Patchell for the stimulating talk there and a useful comment on this paper.
+Patchell for the stimulating talk at MFO and a useful comment on this paper and E.
+Breuillard and I.
+Vigdorovich for insights into $\mathrm{PSL}(d,\mathbb{R})$.
 2. Trees and covariant representations
@@
 See Section 5.2 and Appendix E in [BO] for more on this theme.
-Here we consider a countable tree $\mathrm{T}$ of infinite degree.
+Here we consider a countable tree $\mathrm{T}$ (of infinite degree).
 The tree $\mathrm{T}$ is identified with its vertex set.
@@
 For $t\in\mathrm{T}$ and $\omega\in\partial\mathrm{T}$, we define $f(t,\omega)\coloneq f(t,\omega(n))$ for any $n$ large enough.
-For each $t\in\mathrm{T}$, we consider a copy $\mathrm{T}_{t}$ of $\mathrm{T}$ and equip $\mathrm{T}_{t}$ a compact topology by taking the one-point compactification $\mathrm{T}\cup\{\infty\}$ of $\mathrm{T}$ and identify the new point $\infty$ with $t$.
-That is, every sequence in $\mathrm{T}$ that diverges to infinity converges to $t$ in $\mathrm{T}_{t}$.
+For each $t\in\mathrm{T}$, we consider a copy $\mathrm{T}_{t}$ of $\mathrm{T}$ (or the union of $t$ and its adjacent points) and equip $\mathrm{T}_{t}$ a compact topology by taking the one-point compactification $\mathrm{T}\cup\{\infty\}$ of $\mathrm{T}$ and identify the new point $\infty$ with $t$.
 We consider the embedding
@@
 and equip $\bar{\mathrm{T}}$ with the induced topology.
-This makes $\bar{\mathrm{T}}$ a second-countable compact topological space (Proposition 5.2.5 in [BO]).
+This makes $\bar{\mathrm{T}}$ a second-countable compact topological space such that every automorphism on $\mathrm{T}$ extends to a homeomorphism on $\bar{\mathrm{T}}$ (Proposition 5.2.5 in [BO]).
 Note that if $s\in\mathrm{T}$ and $t_{n}\in\mathrm{T}$ are such that $f(s,t_{n})\to\infty$ in $\mathrm{T}$, then $t_{n}\to s$ in $\bar{\mathrm{T}}$.
 Theorem 4.
-Let $\Lambda$ be a countable discrete group acting on a countable tree $\mathrm{T}$ and on a unital $\mathrm{C}^{*}$-algebra $A$.
-Let $\pi\colon\Lambda\ltimes_{\max}A\to\mathbb{B}(\mathcal{H})$ be a covariant representation.
-Assume that there is a $\Lambda$-equivariant unital positive map from $C(\bar{\mathrm{T}})$ into the center of $A$ and that for every $t]\in\mathrm{T}/\Lambda$, the restriction of $\pi$ to the stabilizer subgroup $\Lambda_{t}\coloneq\{g\in\Lambda:gt=t\}$ is continuous on the reduced crossed product $\Lambda_{t}\ltimes_{\min}A$.
-Then $\pi$ is continuous on the reduced crossed product $\Lambda\ltimes_{\min}A$
+Let $\Lambda$ be a countable discrete group acting on a countable tree $\mathrm{T}$ and on a $\mathrm{C}^{*}$-algebra $C$.
+Let $\pi\colon\Lambda\ltimes_{\max}C\to\mathbb{B}(\mathcal{H})$ be a covariant representation and assume that there is a $\Lambda$-equivariant representation of $C(\bar{\mathrm{T}})$ into the commutant $\pi(C)^{\prime}$ of $\pi(C)$.
+Then $\pi$ is continuous on the reduced crossed product $\Lambda\ltimes_{\min}C$ if for every $t]\in\mathrm{T}/\Lambda$ the restriction of $\pi$ to the stabilizer subgroup $\Lambda_{t}\coloneq\{g\in\Lambda:gt=t\}$ is continuous on the reduced crossed product $\Lambda_{t}\ltimes_{\min}C$.
 Proof.
-Let $\pi\colon\Lambda\ltimes_{\max}A\to\mathbb{B}(\mathcal{H})$ be given.
-We denote by $\lambda$ the unitary representation of $\Lambda$ on $\ell_{2}\mathrm{T}$ induced by $\Lambda\curvearrowright\mathrm{T}$ and denote by $\sigma$ the $\Lambda$ action on $A$.
-We claim that $\pi$ is weakly contained in the representation $\tilde{\pi}$ on $\ell_{2}\mathrm{T}\otimes\mathcal{H}$ given by $g\mapsto(\lambda\otimes\pi)(g)$ for $g\in\Lambda$ and $a\mapsto 1\otimes\pi(a)$ for $a\in A$.
+Let $\pi\colon\Lambda\ltimes_{\max}C\to\mathbb{B}(\mathcal{H})$ be given.
+We denote by $\tau$ the unitary representation of $\Lambda$ on $\ell_{2}\mathrm{T}$ induced by $\Lambda\curvearrowright\mathrm{T}$ and denote by $\sigma$ the $\Lambda$ action on $G$.
+We claim that $\pi$ is weakly contained in the representation $\tilde{\pi}$ on $\ell_{2}\mathrm{T}\otimes\mathcal{H}$ given by $g\mapsto(\tau\otimes\pi)(g)$ for $g\in\Lambda$ and $a\mapsto 1\otimes\pi(a)$ for $a\in C$.
 Indeed, there is a sequence $\zeta_{n}\colon\bar{\mathrm{T}}\to\mathrm{Prob}(\mathrm{T})$ of Borel maps such that $\sup_{x\in\bar{\mathrm{T}}}\|g\zeta_{n}^{x}-\zeta_{n}^{gx}\|_{1}\to 0$ for every $g\in\Lambda$ (Lemma 5.2.6 in [BO]).
@@
 NB: the above sum is an infinite sum of positive Borel functions.
-Let $\theta$ denote the $\Lambda$-equivariant unital completely positive map from $C(\bar{\mathrm{T}})$ into the center of $A$.
-The map $\pi\circ\theta$ canonically extends to a normal unital positive map $\Theta$ from the $\mathrm{C}^{*}$-algebra $B(\bar{\mathrm{T}})$ of the bounded Borel functions into $(\pi\circ\theta)(C(\bar{\mathrm{T}}))^{\prime\prime}$.
-By “normal”, we mean that it sends a bounded pointwise convergent sequence to a strong operator topology convergent sequence.
-We note that $\Theta$ satisfies $|\Theta(f)|\leq\Theta(|f|)$ because the range of $\Theta$ is commutative.
-We define isometries $V_{n}\colon\mathcal{H}\to\ell_{2}\mathrm{T}\otimes\mathcal{H}$ by $V_{n}\xi\coloneq\sum_{t\in\mathrm{T}}\delta_{t}\otimes\Theta(h_{n}^{t})^{1/2}\xi$, or $V_{n}$ is the column operator $(\Theta(h_{n}^{t})^{1/2})_{t\in\mathrm{T}}$.
+The $\Lambda$-equivariant representation of $C(\bar{\mathrm{T}})$ into $\pi(C)^{\prime}$ uniquely extends to a $\sigma$-normal $\Lambda$-equivariant representation $\theta$ from the $\mathrm{C}^{*}$-algebra $B(\bar{\mathrm{T}})$ of bounded Borel functions into $\pi(C)^{\prime}$.
+By $\sigma$-normal”, we mean that it sends a bounded pointwise convergent sequence to a strong operator topology convergent sequence.
+We define isometries $V_{n}\colon\mathcal{H}\to\ell_{2}\mathrm{T}\otimes\mathcal{H}$ by $V_{n}\xi\coloneq\sum_{t\in\mathrm{T}}\delta_{t}\otimes\theta(h_{n}^{t})^{1/2}\xi$, or equivalently, $V_{n}$ is the column operator $(\theta(h_{n}^{t})^{1/2})_{t\in\mathrm{T}}$.
 Then
-$\|V_{n}\pi(g)-(\lambda\otimes\pi)(g)V_{n}\|^{2}=\|\sum_{t}|\Theta(h_{n}^{gt})^{1/2}-\Theta(\sigma_{g}(h_{n}^{t}))^{1/2}|^{2}\|_{\infty}\to 0$
-for every $g\in\Lambda$ and $V_{n}\pi(a)=(1\otimes\pi(a))V_{n}$ for $a\in A$.
+$\|V_{n}\pi(g)-(\tau\otimes\pi)(g)V_{n}\|^{2}=\|\sum_{t}|\theta(h_{n}^{gt})^{1/2}-\theta(\sigma_{g}(h_{n}^{t}))^{1/2}|^{2}\|_{\infty}\to 0$
+for every $g\in\Lambda$ and $V_{n}\pi(a)=(1\otimes\pi(a))V_{n}$ for $a\in C$.
 This proves the claim that $\pi$ is weakly contained in $\tilde{\pi}$.
@@
 For the continuity of $\tilde{\pi}$, it suffices to consider the continuity of the vector states $\varphi$ associated with the vectors of the form $\delta_{t}\otimes\xi$, since they form a cyclic family.
-But $\varphi=\psi\circ E$, where $E$ is the conditional expectation from $\Lambda\ltimes_{\max}A$ onto $\Lambda_{t}\ltimes_{\max}A$ and $\psi$ is the vector state associated with $\xi$.
-(We do not need the fact that the natural embedding $\Lambda_{t}\ltimes_{\max}A\subset\Lambda\ltimes_{\max}A$ is faithful.) Hence the continuity on the reduced crossed product $\Lambda\ltimes_{\min}A$ follows from that for $\Lambda_{t}\ltimes_{\min}A$ for $t\in\mathrm{T}$ one from each $\Lambda$-orbit.
+But $\varphi=\psi\circ E$, where $E$ is the canonical conditional expectation from $\Lambda\ltimes_{\max}C$ onto $\Lambda_{t}\ltimes_{\max}C$ and $\psi$ is the vector state associated with $\xi$.
+(We do not need the fact that the natural embedding $\Lambda_{t}\ltimes_{\max}C\subset\Lambda\ltimes_{\max}C$ is faithful.) Hence the continuity on the reduced crossed product $\Lambda\ltimes_{\min}C$ follows from that for $\Lambda_{t}\ltimes_{\min}C$ for $t\in\mathrm{T}$ one from each $\Lambda$-orbit.
 ∎
-3. A space with “tree structure”
-Let $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. By the normal form of $g\in\tilde{\Gamma}$, we mean the expression
-$g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots z^{\epsilon_{l}}a_{l+1}$
-with $a_{i}\in\Gamma$ and $\epsilon_{i}\in\{\pm\}$, which is reduced in the sense that if $i>1$ and $a_{i}=1$, then $\epsilon_{i-1}=\epsilon_{i}$.
-We introduce the tree $\mathrm{T}\coloneq\tilde{\Gamma}/\Gamma$ on which $\tilde{\Gamma}$ acts.
+3. Extreme boundaries and “tree-graded” spaces
+Let $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. By the reduced form of $g\in\tilde{\Gamma}$, we mean the unique expression
+$g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots a_{l}z^{\epsilon_{l}}a_{l+1}$
+with $a_{i}\in\Gamma$ and $\epsilon_{i}\in\{\pm\}$ such that $i>1$ and $a_{i}=1$ implies $\epsilon_{i-1}=\epsilon_{i}$.
+We introduce the tree structure on $\mathrm{T}\coloneq\tilde{\Gamma}/\Gamma$ on which $\tilde{\Gamma}$ acts.
 We declare that the points $s,t\in\mathrm{T}$ are adjacent if $s^{-1}t\in\{\Gamma z^{\pm}\Gamma\}$.
@@
 We call the point $o\coloneq\Gamma\in\mathrm{T}$ the origin.
-Every $t\in\mathrm{T}$ has unique representative $\rho(t)\in\tilde{\Gamma}$ whose normal form ends by $z^{\pm}$, or $\rho(o)=1$ if $t=o$.
+Every $t\in\mathrm{T}$ has unique representative $\rho(t)\in\tilde{\Gamma}$ whose reduced form ends by $z^{\pm}$, or $\rho(o)=1$ if $t=o$.
 Note that $\rho(at)=a\rho(t)$ for all $a\in\Gamma$ and $t\in\mathrm{T}\setminus\{o\}$ and that $\rho(z^{\pm}t)=z^{\pm}\rho(t)$ for all $t\in\mathrm{T}$.
 Suppose first that $s$ and $t$ are adjacent.
-We label the oriented edge from $t$ to $s$ by $\ell(t,s)\coloneq\kappa(\rho(t)^{-1}s)\in\Gamma\{z_{\pm}\}$, where $\kappa(az^{\pm}\Gamma)\coloneq az_{\pm}$.
-Note that $t$ and $\ell(t,s)$ uniquely determines $s$.
-E.g., $\ell(az\Gamma,azbz\Gamma)=bz_{+}$ and $\ell(azbz\Gamma,az\Gamma)=z_{-}$ for every $a,b\in\Gamma$.
+We label the oriented edge from $t$ to $s$ by $\ell(t,s)\coloneq\kappa(\rho(t)^{-1}s)\in\Gamma\{z^{\pm}\}$, where $\kappa(az^{\pm}\Gamma)\coloneq az^{\pm}$.
+Note that $t$ and $\ell(t,s)$ together uniquely determines $s$.
+E.g., $\ell(az\Gamma,azbz\Gamma)=bz$ and $\ell(azbz\Gamma,az\Gamma)=z^{-1}$ for every $a,b\in\Gamma$.
 The Figure 1 shows how $\mathrm{T}$ and $\ell$ look like.
-Next we extend the label on $(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$ by setting $\ell(t,s)\coloneq\ell(t,t_{1})$, where $t_{1}$ is the unique point between $t$ and $s$ that is adjacent to $t$, or colloquially, $\ell(t,s)$ is the first $\Gamma\{z^{\pm}\}$ component in the normal form $\rho(t)^{-1}\rho(s)$.
-For example, $\ell(az,azbzcz\Gamma)=bz_{+}$ and $\ell(cz\Gamma,azbz\Gamma)=z_{-}$ unless $a=c$.
+Next we extend the label map on $(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$ by setting $\ell(t,s)\coloneq\ell(t,t_{1})$, where $t_{1}$ is the unique point between $t$ and $s$ that is adjacent to $t$, or colloquially, $\ell(t,s)$ is the first $\Gamma\{z^{\pm}\}$ component in the reduced form of $\rho(t)^{-1}\rho(s)$.
+For example, $\ell(az,azbzcz\Gamma)=bz$ and $\ell(cz\Gamma,azbz\Gamma)=z^{-1}$ unless $a=c$.
 Figure 1. The tree $\mathrm{T}$
@@
 Lemma 5.
-Let $(t,s)\in(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$.
-For $a\in\Gamma$, one has $\ell(at,as)=\ell(t,s)$ for $t\neq o$ and $\ell(o,as)=a\ell(o,s)$.
-Also, $\ell(zt,zs)=\ell(t,s)$.
-Let $\Gamma\curvearrowright X$ be a topologically-free extreme boundary with an axial sequence $(z_{n})_{n}$.
-We denote by $\varphi_{n}\colon\tilde{\Gamma}\to\Gamma$ the homomorphism that is identity on $\Gamma$ and send $z$ to $z_{n}\in\Gamma$.
+Let $(t,s)\in(\mathrm{T}\times\bar{\mathrm{T}})$ be a non-diagonal point. For $a\in\Gamma$, one has $\ell(o,as)=a\ell(o,s)$ and $\ell(at,as)=\ell(t,s)$ for $t\neq o$. Also, $\ell(zt,zs)=\ell(t,s)$.
+Let $\Gamma\curvearrowright X$ be an extreme boundary with an axial sequence $(z_{n})_{n}$.
+We denote by $\varphi_{n}\colon\tilde{\Gamma}\to\Gamma$ the homomorphism that is identity on $\Gamma$ and sends $z$ to $z_{n}\in\Gamma$.
+A simple ping-pong argument on $X$ shows that $(\varphi_{n})_{n}$ is asymptotically injective and hence $\Gamma$ is mixed-identity-free.
+More precisely, the following is true.
 Lemma 6.
-Let $g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots z^{\epsilon_{l}}$ be an element in the normal form with $l\geq 1$ and $x\in X\setminus\{z_{\bar{\epsilon}_{l}}\}$.
+Let $g=a_{1}z^{\epsilon_{1}}\cdots z^{\epsilon_{l}}a_{l+1}$ be an element in the reduced form with $l\geq 1$ and $x\in X\setminus\{a_{l+1}^{-1}z_{\bar{\epsilon}_{l}}\}$, where $\bar{\epsilon}$ is the opposite of $\epsilon$.
 Then $\varphi_{n}(g)x\to a_{1}z_{\epsilon_{1}}$ as $n\to\infty$.
-Proof.
-Let $U_{\pm}$ be arbitrarily small neighborhoods of $z_{\pm}$ that satisfy $x\notin U_{\bar{\epsilon}_{l}}$ and $a_{i}(U_{+}\cup U_{-})\cap(U_{+}\cup U_{-})=\emptyset$.
-Then by the ping-pong argument, one has $\varphi_{n}(g)x\in a_{1}U_{\epsilon_{1}}$ for all $n$ large enough.
-∎
-We denote by $\sigma$ the corresponding action of $\Gamma$ on $C(X)\subset C(X)^{\mathcal{U}}$, where the embedding is the constant embedding.
+We denote by $\sigma$ the corresponding action of $\Gamma$ on $C(X)\subset C(X)^{\mathcal{U}}$, where the embedding is the diagonal embedding.
 We denote by $\sigma_{z}$ the automorphism arising from the sequence $(\sigma_{z_{n}})_{n}$.
-They together give rise to the action of $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$ on $C(X)^{\mathcal{U}}$, which is given by
-$\sigma_{g}([f_{n}]_{n})=[\sigma_{\varphi_{n}(g)}(f_{n})]_{n}.$
-We are interested in the commutative $\mathrm{C}^{*}$-algebra
-$C(X_{\mathrm{T}})\coloneq\mathrm{C}^{*}(\bigcup_{g\in\tilde{\Gamma}}\sigma_{g}(C(X)))\subset C(X)^{\mathcal{U}}$
-on which $\tilde{\Gamma}$ acts, where $X_{\mathrm{T}}$ is the Gelfand spectrum of $C(X_{\mathrm{T}})$. By definition, there is a surjective homomorphism
+They together give rise to the action of $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$ on $C(X)^{\mathcal{U}}$, which is given by $\sigma_{g}([f_{n}]_{n})=[\sigma_{\varphi_{n}(g)}(f_{n})]_{n}$.
+We are interested in the $\tilde{\Gamma}-$invariant $\mathrm{C}^{*}$-subalgebra
+$C(X_{\mathrm{T}})\coloneq\mathrm{C}^{*}(\bigcup_{g\in\tilde{\Gamma}}\sigma_{g}(C(X)))\subset C(X)^{\mathcal{U}},$
+where $X_{\mathrm{T}}$ is the Gelfand spectrum of $C(X_{\mathrm{T}})$. By definition, there is a surjective homomorphism
 $Q\colon C(X^{\mathrm{T}})\cong\bigotimes_{t\in\mathrm{T}}C(X)\to C(X_{\mathrm{T}}),\quad Q(f^{(s)})=\sigma_{\rho(s)}(f),$
-where $f^{(s)}\in C(X^{\mathrm{T}})$ is defined for $f\in C(X)$ and $\mathbf{x}=(x_{t})_{t}$ by $f^{(s)}(\mathbf{x})\coloneq f(x_{s})$.
+where $f^{(s)}\in C(X^{\mathrm{T}})$ is defined for $f\in C(X)$ by $f^{(s)}(\mathbf{x})\coloneq f(x_{s})$, $\mathbf{x}=(x_{t})_{t}$.
 This gives rise to an embedding $Q_{*}\colon X_{T}\to X^{\mathrm{T}}$.
@@
 Thus we view $X_{\mathrm{T}}$ as a compact subset of $X^{\mathrm{T}}$.
-For every $s\in\mathrm{T}$, we define the continuous embedding $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ by $\theta(s,x)_{s}\coloneq x$ and $\theta(s,x)_{t}\coloneq\ell(t,s)$ for $t\neq s$.
+We view the label map $\ell$ as a map taking values in $X$ by identifying $\Gamma\{z^{\pm}\}$ with $\Gamma\{z_{\pm}\}$.
+For each $s\in\mathrm{T}$, we define the continuous embedding $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ by $\theta(s,x)_{s}\coloneq x$ and $\theta(s,x)_{t}\coloneq\ell(t,s)$ for $t\neq s$.
 We also define the continuous embedding $\theta\colon\partial\mathrm{T}\to X^{\mathrm{T}}$ by $\theta(\omega)_{t}\coloneq\ell(t,\omega)$.
-Note that $\omega\to\theta(\omega)_{t}$ is locally constant.
-The following theorem asserts that $\Gamma\curvearrowright X_{\mathrm{T}}$ has a “tree structure”. See Figure 2.
-Figure 2. The space $X_{\mathrm{T}}$ around $\theta(s,X)$, where $s=az\Gamma$ and $t=azbz\Gamma$, $\theta(s,bz_{+})=\theta(t,z_{-})$ and $\theta(s,z_{-})=\theta(o,az)$.
+Note that $\omega\to\theta(\omega)_{t}$ is locally constant for every $t$.
+The space $X_{\mathrm{T}}$ has a kind of “tree-graded” structure ([DS]), as follows. See Figure 2.
+Figure 2. The space $X_{\mathrm{T}}$ around $\theta(az\Gamma,bz_{+})=\theta(azbz\Gamma,z_{-})$.
 Theorem 8.
-In the above setting for $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$, the following hold.
+In the above setting for $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}\curvearrowright\mathrm{T}$, the following hold.
   1. (1)
@@
 If $(t_{n})_{n}$ is converges to some point $s^{\prime}$, then by the “if” part and the compactness of $X$, one has $\theta(t_{n},x_{n})\to\theta(s^{\prime},y)$, where $y$ is the limit of $y_{n}\coloneq x_{n}$ if $t_{n}=s^{\prime}$ and $y_{n}\coloneq\ell(s^{\prime},t_{n})$ if $t_{n}\neq s^{\prime}$.
-Since $\theta(s^{\prime},y)=\theta(s,x)$, one has either $s^{\prime}=s$ or $s^{\prime}$ is an adjacent point to $s$ that satisfies $x=\ell(s,s^{\prime})$ and $y=\ell(s^{\prime},s)$.
+Since $\theta(s^{\prime},y)=\theta(s,x)$, one has by (3) either $s^{\prime}=s$ or $s^{\prime}$ is an adjacent point to $s$ that satisfies $x=\ell(s,s^{\prime})$ and $y=\ell(s^{\prime},s)$.
 This proves the “only if” part for the case $(t_{n})_{n}$ is convergent.
@@
 If $x_{t}\in X\setminus\Gamma\{z_{\pm}\}$, then $w(t)\coloneq t$.
-If $x_{t}\in\Gamma\{z_{\pm}\}$, then $w(t)\coloneq\rho(t)\kappa^{-1}(x_{t})$, where $\kappa^{-1}(az_{\pm})\coloneq az^{\pm}\Gamma\in\mathrm{T}$.
-Note that $w(t)$ is adjacent to $t$ in the latter case.
+If $x_{t}\in\Gamma\{z_{\pm}\}$, then we define $w(t)$ to be the unique point that is adjacent to $t$ and satisfies $\ell(t,w(t))=x_{t}$.
 We claim that if $s,t\in\mathrm{T}$ are adjacent, then it cannot happen that $w(s)\neq t$ and $w(t)\neq s$ simultaneously.
@@
 It follows that $\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})\supset X_{\mathrm{T}}$.
-The assertion (9) is a consequence of (1-8). ∎
+Finally, the assertion (9) is a consequence of (1-8). ∎
 4. Proof of Theorem 1
 Proof of Theorem 1.
-By Theorem 2.6 in [Ro], $\mathrm{C}^{*}$-selflessness follows from the “more precisely” statement.
-Let $\Gamma\curvearrowright X$ be a topologically-free extreme boundary with an axial sequence $(z_{n})_{n}$ and $\pi\colon\Gamma*\mathopen{\langle}z\mathclose{\rangle}\to\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ be the corresponding homomorphism.
-We set $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$.
-We take a $\Gamma$-equivariant embedding $C(X)\subset\ell_{\infty}\Gamma\subset\mathbb{B}(\ell_{2}\Gamma)$.
-It gives rise to $\tilde{\Gamma}-$equivariant embedding of $C(X_{\mathrm{T}})$ into $\mathbb{B}(\ell_{2}\Gamma)^{\mathcal{U}}$, where $C(X_{\mathrm{T}})$ is defined in the previous section.
-We take a faithful representation of $\mathbb{B}(\ell_{2}\Gamma)^{\mathcal{U}}$ on a Hilbert space $\mathcal{H}$ and view $\pi$ as a covariant representation of $\tilde{\Gamma}\ltimes C(X_{\mathrm{T}})$.
-Then the representation $\pi|_{C(X_{\mathrm{T}})}$ canonically extends to a normal representation $\pi$ of the $\mathrm{C}^{*}$-algebra $B(X_{\mathrm{T}})$ of the bounded Borel functions on $X_{\mathrm{T}}$.
-Let $p$ denote the orthogonal projection on $\mathcal{H}$ corresponding to the countable subset $\tilde{\Gamma}\{z_{\pm}\}$ of $X_{\mathrm{T}}$, which is possibly zero.
-The unitary representation $p\pi$ of $\tilde{\Gamma}$ is covariant with $\ell_{\infty}\tilde{\Gamma}$ and hence unitarily equivalent to a multiple of the regular representation.
-By Theorem 6, there is a $\tilde{\Gamma}-$equivariant embedding of $C(\bar{\mathrm{T}})$ into $p^{\perp}\pi(B(X_{\mathrm{T}}))$.
-Thus by Theorem 2 and the fact that $\pi|_{\Gamma}$ is weakly equivalent to the regular representation of $\Gamma$, the unitary representation $p^{\perp}\pi$ of $\tilde{\Gamma}$ is weakly contained in the regular representation.
+[…]
+5. Direct products of $\mathrm{C}^{*}$-selfless groups
+It is not hard to see that if $A\subset B$ is an existential embedding of $\mathrm{C}^{*}$-algebras (see Section 1 in [Ro]), then so is $A\otimes C\subset B\otimes C$, provided that the $\mathrm{C}^{*}$-algebra $C$ is exact.
+(Without the exactness assumption, the canonical embedding $A^{\mathcal{U}}\otimes C\subset(A\otimes C)^{\mathcal{U}}$ is not continuous w.r.t.
+the minimal tensor product.
+See [BO].) Hence, if $A_{i}\subset B_{i}$ are existential embedding and $B_{i}$ are exact, then the corresponding embedding $A_{1}\otimes A_{2}\subset B_{1}\otimes B_{2}$ is existential.
+The same holds true in the $\mathrm{C}^{*}$-probability space setting.
+We denote by $(\mathcal{T},\omega)$ the Toeplitz $\mathrm{C}^{*}$-probability space consisting of the Toeplitz algebra generated by the unilateral shift $T$ on $\ell_{2}(\{0,1\,\ldots\})$ and the vacuum state $\omega$ associated with the vacuum vector $\delta_{0}$.
+We denote by $(\mathcal{C}_{1},\tau)\subset(\mathcal{T},\omega)$ the $\mathrm{C}^{*}$-probability space generated by $s\coloneq(T+T^{*})/2$ in $\mathcal{T}$.
+Note that $s$ corresponds to the identity function under the isomorphism $(\mathcal{C}_{1},\tau)\cong(C([-1,1]),\gamma)$, where the state $\gamma$ is given by by the semicircle probability measure $\frac{2}{\pi}\sqrt{1-s^{2}}\,\mathrm{d}s$.
+See Section 2.6 in [VDN].
+We denote by $\mathbf{F}_{\infty}$ the free group on generators $\{t_{i}\}_{i=1}^{\infty}$.
+We view it as embedded in the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})$ acting on $\ell_{2}\mathbf{F}_{\infty}$.
+It forms the $\mathrm{C}^{*}$-probability space $(\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty}),\tau)$, where $\tau$ is the canonical tracial state that is given by $\tau(g)=\delta_{g,1}$ for $g\in\mathbf{F}_{\infty}$.
+It is well-known ([VDN]) that the sequence
+$s_{n}=(8n)^{-1/2}\sum_{i=1}^{n}(t_{i}+t_{i}^{-1})$
+in $(\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty}),\tau)$ strongly converges to $s$ in $(\mathcal{C}_{1},\tau)$.
+Namely, for every polynomial $p$, one has $\tau(p(s_{n}))\to\gamma(p(s))$ and $\|p(s_{n})\|\to\|p(s)\|$.
+We will prove a vectorial version of this fact in Theorem 9.
+Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space on which $\mathbf{F}_{\infty}$ acts (leaving $\varphi$ invariant).
+Then the reduced crossed product $A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty}$ is equipped with the canonical state, still denoted by $\varphi$, that is given by $\varphi(ag)=\varphi(a)\tau(g)$ for $a\in A$ and $g\in\mathbf{F}_{\infty}$.
+Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space.
+We view $A$ as embedded in the ultrapower $(A,\varphi)^{\mathcal{U}}\coloneq(A^{\mathcal{U}},\varphi^{\mathcal{U}})$ diagonally.
+(The difference in the free ultrafilter $\mathcal{U}$ will have no effect on the discussion.) Consider a bounded sequence $(a_{n})_{n=1}^{\infty}$ of self-adjoint elements in $(A,\varphi)$ that is strongly convergent to $a$ in $(\mathrm{C}^{*}(\{1,a\}),\psi)$.
+We say the sequence $(a_{n})_{n=1}^{\infty}$ is strongly asymptotically free (w.r.t.
+$\mathcal{U}$) from a $\mathrm{C}^{*}$-subalgebra $A_{0}$ if the induced homomorphism
+$(A_{0},\varphi|_{A_{0}})\ast(\mathrm{C}^{*}(\{1,a\}),\psi)\to(A^{\mathcal{U}},\varphi^{\mathcal{U}})$
+is well-defined and continuous on the reduced free product. Note that the reduced free product is compatible with embedding (see [BD] or Corollary 4.8.4 in [BO].
+Theorem 9.
+Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space on which $\mathbf{F}_{\infty}$ acts by $\sigma$. Assume that
+$\forall a\in A\quad\frac{1}{2n}\sum_{i=1}^{n}(\sigma_{t_{i}}(a)+\sigma_{t_{i}}^{-1}(a))\to\varphi(a).$
+Then, the sequence $(s_{n})_{n=1}^{\infty}$ as above in $A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty}$ is strongly asymptotically free from $A$ (in fact from $A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty}$).
+Proof.
+The proof is inspired from [VDN] and [HP]. We can write $\mathbf{F}_{\infty}$ as a disjoint union
+$\mathbf{F}_{\infty}=\{1\}\cup\bigcup_{i}(E_{i}^{+}\cup E_{i}^{-}),$
+where $E_{i}^{\pm}$ is the set of those elements whose reduced form starts with (a positive power of) $t_{i}^{\pm}$.
+We denote by $P_{i,\pm}$ the orthogonal projection on $\ell_{2}\mathbf{F}_{\infty}$ that corresponds to $E_{i}^{\pm}$.
+They are mutually orthogonal.
+One has $t_{i}=P_{i,+}t_{i}+t_{i}P_{i,-}\eqcolon u_{i,+}+u_{i,-}^{*}$ (see Proof of Proposition 1.1 in [HP]).
+Note that $u_{i,\pm}$ are partial isometries such that $u_{i,\pm}^{*}u_{i,\pm}=1-P_{i,\mp}$ and $u_{i,\pm}u_{i,\pm}^{*}=P_{i,\pm}$.
+Hence, for
+$T_{n}\coloneq(2n)^{-1/2}\sum_{i=1}^{n}(u_{i,+}+u_{i,-}),$
+one has $s_{n}=(T_{n}+T_{n}^{*})/2$ and $T_{n}^{*}T_{n}=1-(2n)^{-1}\sum_{i=1}^{n}(P_{i,-}+P_{i,+})\approx 1$.
+We represent $A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty}$ on the Hilbert space $\mathcal{H}\coloneq L^{2}(A,\varphi)\otimes\ell_{2}\mathbf{F}_{\infty}$ by $g\mapsto(1\otimes g)$ for $g\in\mathbf{F}_{\infty}$ and $a\mapsto\pi(a)$, $\pi(a)(\xi\otimes\delta_{h})=\sigma_{h}^{-1}(a)\xi\otimes\delta_{h}$, for $a\in A$.
+Then, by the assumption,
+$(1\otimes T_{n})^{*}\pi(a)(1\otimes T_{n})=\pi(\frac{1}{2n}\sum_{i=1}^{n}(\sigma_{g_{i}}^{-1}(a)(1-P_{i,-})+\sigma_{g_{i}}(a)(1-P_{i,+}))\to\varphi(a)$
+for $a\in A$.
+Hence $T\coloneq[(1\otimes T_{n})]_{n}\in\mathbb{B}(\mathcal{H})^{\mathcal{U}}$ is an isometry that satisfies $T^{*}\pi(a)T=\varphi(a)$ for $a\in A$.
+We denote by $\omega$ the vector state on $\mathbb{B}(\mathcal{H})$ associated with the vector $\hat{1}\otimes\delta_{1}$.
+Thus $(A,\varphi)\subset(\mathbb{B}(\mathcal{H}),\omega)$.
+The isometry $T$ in $(\mathbb{B}(\mathcal{H}),\omega)^{\mathcal{U}}$ generates a copy of the Toeplitz probability space $(\mathcal{T},\omega)$.
+In particular, $s_{n}]_{n}=(T+T^{*})/2$ generates a copy of $(\mathcal{C}_{1},\tau)$ in $(A\rtimes\mathbf{F}_{\infty},\varphi)^{\mathcal{U}}$.
+It is rather easy to see that $A$ and $\mathrm{C}^{*}(T)$ are free from each other w.r.t.
+$\omega^{\mathcal{U}}$.
+It remains to show continuity of the induced homomorphism from the reduced free product $(A,\varphi)\ast(\mathcal{T},\omega)$ into $(\mathbb{B}(\mathcal{H}),\omega)^{\mathcal{U}}$.
+The universal $\mathrm{C}^{*}$-algebra generated by $A$ and an isometry $T$ that satisfy $T^{*}aT=\varphi(a)$ is the Toeplitz–Pimsner algebra $\mathcal{T}_{A,\varphi}$ over the Hilbert $A$-module $\mathop{\overline{\mathrm{span}}}ATA$ (see, e.g., Example 4.6.11 in [BO]).
+Since the canonical conditional expectation from $\mathcal{T}_{A,\varphi}$ onto $A$ is non-degenerate (Theorem 4.6.6 in [BO]), the Toeplitz–Pimsner algebra $\mathcal{T}_{A,\varphi}$ is isomorphic to the reduced free product $(A,\varphi)*(\mathcal{T},\omega)$.
+This proves the desired continuity.
 ∎
-A crossed product analogue of Theorem 1 holds true under the additional assumption that $\Gamma$ is exact, which ensures $\Gamma\ltimes_{\min}A^{\mathcal{U}}\subset(\Gamma\ltimes_{\min}A)^{\mathcal{U}}$.
+Proof of Theorem 2.
+By Theorem 4.1 in [Ro], we may assume that $i\in\{1,2\}$. Since $(A_{i},\varphi_{i})$ are selfless, the embeddings
+[…]
+are existential (Theorem 2.6 in [Ro]).
+The $\mathrm{C}^{*}$-algebras $B_{i}$ are exact by Dykema’s theorem ([Dy] or Corollary 4.8.3 in [BO]).
+Hence the embedding of $(A_{1}\otimes A_{2},\varphi_{1}\otimes\varphi_{2})$ into $(B_{1}\otimes B_{2},\psi_{1}\otimes\psi_{2})$ is existential.
+We write $\mathbf{F}_{\infty}^{(i)}$ for the copy of $\mathbf{F}_{\infty}$ in $B_{i}$.
+Note that
+[…]
+$\supset(A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty},\varphi),$
+[…]
+$\|\sum_{j=1}^{m}x_{j}\otimes y_{j}\|\leq 3(\sum_{j=1}^{m}\|x_{j}\|^{2}\|y_{j}\|^{2})^{1/2}\precsim m^{1/2}\max\|x_{j}\|\|y_{j}\|$
+whenever $x_{j}$’s (or $y_{j}$’s) are freely independent and mean zero. ∎
+6. Completely selfless $\mathrm{C}^{*}$-algebras
+In this section, we indicate the way to circumvent the exactness assumption in Theorem 2.
+Recall that an embedding $A\subset B$ of $\mathrm{C}^{*}$-algebras is said to be existential if there is an ultrafilter $\mathcal{U}$ and an embedding $\sigma\colon B\hookrightarrow A^{\mathcal{U}}$ whose restriction to $A$ is the diagonal embedding of $A$ into $A^{\mathcal{U}}$.
+We say that the embedding is completely existential if there is $\sigma$ as above that moreover satisfies that for every $\mathrm{C}^{*}$-algebra $C$ the induced homomorphism $B\otimes C\to(A\otimes C)^{\mathcal{U}}$ is continuous.
+By the Effros–Haagerup lifting theorem, the additional condition is equivalent to that the homomorphism $\sigma$ is locally liftable to completely positive contractions into the $\ell_{\infty}$-sum $\prod A$.
+Namely, $A\subset B$ is completely existential if and only if for every finite dimensional operator system $E\subset B$ and $\varepsilon>0$, there is a unital completely positive map $\theta\colon E\to A$ satisfying that $\|\theta(x)\|\geq(1-\varepsilon)\|x\|$ for $x\in E$, $\|\theta(xy)-\theta(x)\theta(y)\|<\varepsilon$ for those $x,y\in E$ such that $xy\in E$, and $\|\theta(a)-a\|<\varepsilon$ for $a\in E\cap A$.
+In the $\mathrm{C}^{*}$-probability space setting $(A,\varphi)\subset(B,\psi)$, it is moreover required that $\|\varphi\circ\theta-\psi|_{E}\|<\varepsilon$.
+It is not hard to see that the class of completely existential embeddings are closed under compositions and tensor products.
+It is also closed under reduced free products (cf.
+Corollary 1.9 in [Ro]).
+Lemma 10.
+For every completely existential embedding $(A,\varphi)\subset(B,\psi)$ of $\mathrm{C}^{*}$-probability spaces and every $(C,\rho)$, the embedding $(A,\varphi)\ast(C,\rho)\subset(B,\psi)\ast(C,\rho)$ is completely existential.
+Proof.
+Wthout loss of generality, we assume that $B$ is separable.
+Take an increasing sequence of finite dimensional operator systems $E_{n}\subset B$ such that $\bigcup E_{n}$ is dense in $B$ and $\bigcup(E_{n}\cap A)$ is dense in $A$.
+Since the embedding $(A,\varphi)\subset(B,\psi)$ is completely existential, for each $n$, there is a unital completely positive map $\theta_{n}\colon E_{n}\to A$ that is approximately isometric, approximately multiplicative, approximately state preserving, and $\theta_{n}|_{E_{n}\cap A}\approx\mathrm{id}_{E_{n}\cap A}$.
+It extends to a unital completely positive map, still denoted by $\theta_{n}$, from $B$ into $\mathbb{B}(L^{2}(A,\varphi))$.
+We denote by $\omega$ the vector state that corresponds to the GNS-vector $\hat{1}$ in $L^{2}(A,\varphi)$.
+One has $\psi_{n}\coloneq\omega\circ\theta_{n}\to\psi$ pointwise on $B$.
+Let $B_{n}$ denote the quotient of $B$ by the GNS-kernel of $\psi_{n}$.
+By the Choda–Blanchard–Dykema theorem ([BD] or Theorem 4.8.5 in [BO]), $\theta_{n}\ast\mathrm{id}_{C}$ defines a unital completely positive map from $(B_{n},\psi_{n})\ast(C,\rho)$ into $(\mathbb{B}(L^{2}(A,\varphi)),\omega)\ast(C,\rho)$.
+By the Skoufranis–Pisier theorem (Theorem 7.1 in [Pi]) for amalgamated free products, the (a priori formally defined) embedding
+$(B,\psi)\ast(C,\rho)\hookrightarrow\prod_{n\in\mathbb{N}}(B_{n},\psi_{n})\ast(C,\rho)/\mathcal{U}$
+is continuous and locally liftable.
+By composing this with $\theta_{n}\ast\mathrm{id}_{C}]_{n}$, one obtains a locally liftable embedding of $(B,\psi)\ast(C,\rho)$ into $((\mathbb{B}(L^{2}(A,\varphi)),\omega)\ast(C,\rho))^{\mathcal{U}}$, which lands in $((A,\varphi)\ast(C,\rho))^{\mathcal{U}}$.
+This witnesses existentialness.
+∎
+We say a $\mathrm{C}^{*}$-probability space completely selfless if the “complete” analogue of the equivalent conditions in Theorem 2.6 in [Ro] hold.
+A group is completely $\mathrm{C}^{*}$-selfless if its reduced group $\mathrm{C}^{*}$-algebra is completely selfless.
+Note that no nuclear tracial $\mathrm{C}^{*}$-probability space is completely selfless, since if it were, then the tracial state on $\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})$ would be amenable (see Chapter 6 in [BO]), which is absurd.
+Theorems 1 and 2 are upgraded as follows.
+Theorem 11.
+An infinite countable discrete group having a topologically-free extreme boundary is completely $\mathrm{C}^{*}$-selfless.
+The tensor product of separable and completely selfless $\mathrm{C}^{*}$-probability spaces is completely selfless.
+In particular, the class of countable discrete groups that are completely $\mathrm{C}^{*}$-selfless is closed under direct product.
+The proof is same as Theorems 1 and 2, but uses the following fact ([Sh] or Example 4.6.11 and Exercise 4.8.1 in [BO]).
+Lemma 12.
+Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space, $(\mathcal{T},\omega)$ be the Toeplitz $\mathrm{C}^{*}$-probability space, and $C$ be a unital $\mathrm{C}^{*}$-algebra.
+Then $C\otimes((A,\varphi)\ast(\mathcal{T},\omega))$ is the universal $\mathrm{C}^{*}$-algebra generated by $C\otimes A$ and an isometry $T$ (the generator of the Toeplitz algebra) that satisfies $T^{*}(c\otimes a)T=\varphi(a)(c\otimes 1)$ for $a\in A$ and $c\in C$.
+Moreover, the free product state $\psi$ on $(A,\varphi)\ast(\mathcal{T},\omega)$ is the unique state that satisfies $\psi|_{A}=\varphi$ and $\psi(aTT^{*}a^{*})=0$ for $a\in A$.
+We do not use it, but note the fact that $(A,\varphi)\ast(\mathcal{T},\omega)$ is simple and purely infinite, provided that $A\cap\mathbb{K}(L^{2}(A,\varphi))=0$ ([Ku]).
+Theorem 13.
+Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space.
+Suppose that there are a $\mathrm{C}^{*}$-probability space $(B,\psi)$ containing $(A,\varphi)$, an ultrafilter $\mathcal{U}$, and an isometry $T\in B^{\mathcal{U}}$ that satisfies that $T^{*}aT=\varphi(a)$ and $\psi^{\mathcal{U}}(aTT^{*}a^{*})=0$ for all $a\in A$ and that $(T+T^{*})/2\in A^{\mathcal{U}}$.
+Then $(A,\varphi)$ is completely selfless.
+Proof.
+Let $(A_{1},\varphi_{1})\subset(A_{2},\varphi_{2})$ denote $(A,\varphi)*(\mathcal{C}_{1},\tau)\subset(A,\varphi)*(\mathcal{T},\omega)$, see Section 5.
+Let a $\mathrm{C}^{*}$-probability space $(C,\theta)$ be given.
+We have to show there is an embedding
+$(C\otimes A_{1},\theta\otimes\varphi_{1})\hookrightarrow((C\otimes A)^{\mathcal{U}},(\theta\otimes\varphi)^{\mathcal{U}})$
+that extends the diagonal embedding of $C\otimes A$.
+We view $T$ as $1_{C}\otimes T$ in $(C\otimes B)^{\mathcal{U}}$.
+Then by Lemma 12, the isometry $T$ gives rise to a representation of $C\otimes A_{2}$ in $(C\otimes B)^{\mathcal{U}}$.
+Since the state $(\theta\otimes\psi)^{\mathcal{U}}$ on $(C\otimes B)^{\mathcal{U}}$ restricts to $\theta\otimes\varphi_{2}$ on $C\otimes A_{2}$, this representation is a state-preserving embedding of $C\otimes A_{2}$ into $(C\otimes B)^{\mathcal{U}}$.
+Since the generator $(T+T^{*})/2$ of $\mathcal{C}_{1}$ belongs to $A^{\mathcal{U}}$, it restricts to a desired embedding of $C\otimes A_{1}$ into $(C\otimes A)^{\mathcal{U}}$.
+∎
+7. Proof of Theorem 3
+Proof of Theorem 3.
+We deal with the first case.
+By Glimm’s lemma, $\varphi$ is approximated on any finite subset $F$ of $A$ by a pure state $\psi$ that are disjoint from $\varphi$.
+By excision (see Proposition 11.4.2 in [BO]), there is a net of projections $p_{j}$ such that $\|p_{j}ap_{j}-\psi(a)p_{j}\|\to 0$ for $a\in A$.
+The net $(p_{j})_{j}$ converges in $A^{**}$ to the support projection of the pure state $\psi$.
+Hence $\varphi(ap_{j}a^{*})\to 0$ for $a\in A$.
+Thus, by pure infiniteness, there are a directed set $I$ and a net $(T_{i})_{i\in I}$ of isometries in $A$ such that $\|T_{i}^{*}aT_{i}-\varphi(a)\|\to 0$ and $\varphi(aT_{i}T_{i}^{*}a^{*})\to 0$ for $a\in A$.
+Let $\mathcal{U}$ be a cofinal ultrafilter on $I$.
+The isometry $T\coloneq[T_{i}]_{i}$ in $A^{\mathcal{U}}$ satisfies $T^{*}aT=\varphi(a)$ and $\varphi^{\mathcal{U}}(aTT^{*}a^{*})=0$ for $a\in A$.
+By Theorem 13, we are done.
+We deal with the second case.
+Since $A$ is $\mathcal{Z}$-stable and $\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})\hookrightarrow\mathcal{Z}^{\mathcal{U}}$ (Theorem 4.1 in [Oz]), the embedding $A\subset A\otimes C^{*}_{\lambda}(\mathbf{F}_{\infty})$ is existential by exactness, where we identify $A$ with $A\otimes\mathbb{C}1$.
+Since $A$ is simple and monotracial, the uniform Dixmier property holds by the Haagerup–Zsido theorem ([HZ]).
+Namely, there is a net $(F_{i})_{i\in I}$ of finite families of unitary elements in $A$ that satisfies $|F_{i}|^{-1}\sum_{u\in F_{i}}uau^{*}\to\tau(a)$ for $a\in A$.
+By replacing $F_{i}$ with $F_{i}^{*}F_{i}$, we may assume that the finite set $F_{i}$ is closed under the adjoint operation.
+Let $\{u^{(i)}_{n}:n=1,\ldots,|F_{i}|\}$ be an enumeration of $F_{i}$.
+Since $A\subset A\otimes\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})$ is existential, by replacing $u^{(i)}_{n}$ with likenesses in $A$ of $u^{(i)}_{n}\otimes t_{n}$ in $A\otimes\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})$, we may further assume by Voiculescu’s inequality (see Proof of Theorem 2) that $\||F_{i}|^{-1}\sum_{u\in F_{i}}uau\|\to 0$ for $a\in A$.
+Let $\mathcal{U}$ be a cofinal ultrafilter on $I$.
+Let’s consider the Cuntz algebra $(\mathcal{O}_{\infty},\omega)$ generated by isometries $\{l_{n}\}$ with mutually orthogonal ranges, together with the vacuum state.
+Recall that $\{(l_{n}+l_{n}^{*})/2\}$ generates the free semicircular system $(\mathcal{C},\tau)\subset(\mathcal{O}_{\infty},\omega)$ (see [VDN]).
+The elements
+$T_{i}\coloneq(2|F_{i}|)^{-1/2}\sum_{n}(u^{(i)}_{n}+(u^{(i)}_{n})^{*})\otimes l_{n}$
+in $A\otimes\mathcal{O}_{\infty}$ satisfies $T_{i}^{*}aT_{i}\to\tau(a)$ and $(\tau\otimes\omega)(aT_{i}T_{i}^{*}a^{*})=0$ for $a\in A$.
+Moreover, $(T_{i}+T_{i}^{*})/2\in A\otimes\mathcal{C}$.
+Hence by Lemma 12, it gives rise to an embedding of $(A,\tau)\ast(\mathcal{T},\omega)$ into $(A\otimes\mathcal{O}_{\infty},\tau\otimes\omega)^{\mathcal{U}}$.
+It restricts to an embedding of $(A,\tau)\ast(\mathcal{C}_{1},\tau)$ into $(A\otimes\mathcal{C},\tau\otimes\tau)^{\mathcal{U}}$ that witnesses the relative existentialness (see [Ro]) of $(A,\tau)\hookrightarrow(A,\tau)\ast(\mathcal{C}_{1},\tau)$ in $(A\otimes\mathcal{C},\tau\otimes\tau)$.
+Since the embedding of $(A,\tau)$ into $(A\otimes\mathcal{C},\tau\otimes\tau)$ is existential by assumption, it follows that the embedding of $(A,\tau)$ into $(A,\tau)\ast(\mathcal{C}_{1},\tau)$ is also existential.
+∎
+8. The Powers property and selflessness
+$\mathrm{C}^{*}$-simplicity of the free group of rank $2$ has been proved by R.
+T.
+Powers ([Po]) by a combinatorial method.
+The combinatorial method is later formalized by P.
+de la Harpe ([dlH]) as the Powers property.
+The following is very close to it in spirit.
+It is designed to allow the Haagerup–Pisier type decomposition ([HP]).
+Definition.
+We say that a group $\Gamma$ has property $\mathrm{P}_{\mathrm{PHP}}$ if for every finite subset $F\subset\Gamma$, $\varepsilon>0$, and $n\in\mathbb{N}$, there are $t_{i}\in\Gamma$ and $C_{i}\subset D_{i}\subset\Gamma$, $i=1,\ldots,n$, satisfying that the members of
+$\{aC_{i}:a\in F,\,i=1,\ldots,n\}\cup\{bt_{j}^{-1}(\Gamma\setminus D_{j}):b\in F,\,j=1,\ldots,n\}$
+are mutually disjoint and that
+$\sup_{x\in\Gamma}|\{i:x\in D_{i}\}\cup\{j:x\in t_{j}^{-1}(\Gamma\setminus C_{j})\}|\leq\varepsilon n^{1/2}.$
+It is not too hard to see that the class of groups with $\mathrm{P}_{\mathrm{PHP}}$ is closed under the direct product; for $t_{i}^{(k)}\in\Gamma^{(k)}$ and $C_{i}^{(k)}\subset D_{i}^{(k)}\subset\Gamma^{(k)}$, $k=1,2$, set $t_{i}\coloneq(t_{i}^{(1)},t_{i}^{(2)})\in\Gamma^{(1)}\times\Gamma^{(2)}$, $C_{i}\coloneq C_{i}^{(1)}\times C_{i}^{(2)}$, and $D_{i}\coloneq(D_{i}^{(1)}\times\Gamma^{(2)})\cup(\Gamma^{(1)}\times D_{i}^{(2)})$.
+Theorem 14.
+A group $\Gamma$ with property $\mathrm{P}_{\mathrm{PHP}}$ is completely $\mathrm{C}^{*}$-selfless.
+Proof.
+Given a finite subset $F\subset\Gamma$, $\varepsilon>0$, and $n\in\mathbb{N}$, we take $t_{i}$ and $C_{i}\subset D_{i}\subset\Gamma$ as in Definition.
+Let $P_{i}\leq Q_{i}$ denote the diagonal projections in $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to the subsets $C_{i}\subset D_{i}$ and set $P_{i}^{\prime}\coloneq t_{i}^{-1}Q_{i}^{\perp}t_{i}$ and $Q_{i}^{\prime}\coloneq t_{i}^{-1}P_{i}^{\perp}t_{i}$.
+Note that $\{P_{i}\}\cup\{P^{\prime}_{j}\}$ are mutually orthogonal.
+Here $\Gamma$ is represented in $\mathbb{B}(\ell_{2}\Gamma)$ by the left regular representation.
+We denote by $\omega$ the vector state on $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to $\delta_{1}$, which satisfies $\omega|_{\mathrm{C}^{*}_{\lambda}(\Gamma)}=\tau$.
+We define $T\in\mathbb{B}(\ell_{2}\Gamma)$ by
+$T\coloneq(2n)^{-1/2}\sum_{i}(P_{i}t_{i}+t_{i}^{-1}Q_{i}^{\perp})=(2n)^{-1/2}\sum_{i}(P_{i}t_{i}+P^{\prime}_{i}t_{i}^{-1}).$
+The element $T$ is an almost isometry as
+$T^{*}T=(2n)^{-1}\sum_{i}(t_{i}^{-1}P_{i}t_{i}+t_{i}P_{i}^{\prime}t_{i}^{-1})=1-(2n)^{-1}\sum_{i}(Q_{i}^{\prime}+Q_{i})\approx_{\varepsilon n^{-1/2}}1.$
+Moreover, $T^{*}aT=0=\tau(a)$ for every $a\in F\setminus\{1\}$. One has by Schur’s test
+$\|\sum_{i}(Q_{i}-P_{i})t_{i}\|$ $=\|\left[\begin{matrix}Q_{1}-P_{1}&\cdots&Q_{n}-P_{n}\end{matrix}\right]\left[\begin{matrix}t_{1}&&\\ &\ddots&\\ &&t_{n}\end{matrix}\right]\left[\begin{matrix}Q^{\prime}_{1}\\ \vdots\\ Q^{\prime}_{n}\end{matrix}\right]\|$
+$\leq\|\sum_{i}Q_{i}-P_{i}\|^{1/2}\|\sum_{i}Q^{\prime}_{i}\|^{1/2}\leq\varepsilon n^{1/2}$
+and hence
+$\|T+T^{*}-(2n)^{-1/2}\sum_{i}(t_{i}+t_{i}^{*})\|\leq 2(2n)^{-1/2}\|\sum_{i}(Q_{i}-P_{i})t_{i}\|\leq 2\varepsilon.$
+Also, for every finite subset of $\Gamma$, the corresponding diagonal projection $R\in\mathbb{B}(\ell_{2}\Gamma)$ satisfies $\|RT\|^{2}\leq(2n)^{-1}\|\sum_{i}P_{i}+P_{i}^{\prime}\|\|\sum_{i}t_{i}^{-1}Rt_{i}+t_{i}Rt_{i}^{-1}\|\leq n^{-1}\operatorname{rank}R$.
+It follows that a suitable ultra-limit $\tilde{T}\in\mathbb{B}(\ell_{2}\Gamma)^{\mathcal{U}}$ of $T$ as above verifies the condition of Theorem 13 for $(\mathrm{C}^{*}_{\lambda}(\Gamma),\tau)\subset(\mathbb{B}(\ell_{2}\Gamma),\omega)$ and hence $(\mathrm{C}^{*}_{\lambda}(\Gamma),\tau)$ is completely selfless.
+∎
+Proposition 15.
+Let $\Gamma$ be a group.
+Assume that there is a continuous action of $\Gamma$ on a Hausdorff topological space that is minimal, extremely proximal, and topologically free.
+Then $\Gamma$ has property $\mathrm{P}_{\mathrm{PHP}}$.
+Proof.
+We sketch the proof by adopting that of Lemma 4 in [dlH].
+Let a finite subset $F\subset\Gamma$ and $n$ be given.
+Then, there are $t_{1},\ldots,t_{n}$ in $\Gamma$ and mutually disjoint open subsets $U_{i}^{\pm}$ such that $t_{i}(L\setminus U_{i}^{-})\subset U_{i}^{+}$ for every $i$.
+Moreover, we may assume that $a(U_{i}^{+}\cup U_{i}^{-})\cap(U_{j}^{+}\cup U_{j}^{-})=\emptyset$ for every $a\in F\setminus\{1\}$ and every $i,j$.
+We fix a point $x_{0}\in L$ and set $C_{i}\coloneq D_{i}\coloneq\{s\in\Gamma:sx_{0}\in U_{i}^{+}\}$.
+Then they witness $\mathrm{P}_{\mathrm{PHP}}$ with $\varepsilon=0$.
+∎
+Acylindrically hyperbolic groups with trivial finite radical satisfy the assumption of Proposition 15 (Proposition 0.3 in [AD]).
+Hence it removes property RD assumption from Amrutam et al.’s theorem ([AGKEP]), see also [Ya] for an alternative proof.
+However, unlike Amrutam et al.
+and Yang’s theorem ([AGKEP] [Ya]), it does not provide an element in $\Gamma^{\mathcal{U}}$ that is strongly free from $\Gamma$.
+The same comment applies to Theorem 1 and the following proposition that partly generalizes Vigdorovich’s theorem ([Vi]).
+Proposition 16.
+Every Zariski-dense subgroup $\Gamma\leq\mathrm{PSL}(d\geq 2,\mathbb{R})$ satisfies property $\mathrm{P}_{\mathrm{PHP}}$.
+Proof.
+We follow the proof of Proposition 1 in [BCH].
+Let $B:=\mathbb{P}^{d-1}(\mathbb{R})$ denote the homogeneous space of $G\coloneq\mathrm{PSL}(d,\mathbb{R})$, on which $\Gamma$ acts continuously.
+Take $t_{0}\in\Gamma$ with eigenvalues $\lambda_{1}>\lambda_{2}>\cdots>\lambda_{n}>0$ and denote by $\xi_{1}:\cdots:\xi_{n}]$ the corresponding homogeneous coordinates on $B$.
+We set $p\coloneq[1:0:\cdots:0]\in K\coloneq\{[\xi_{1}:\cdots:\xi_{n}]:\xi_{n}=0\}$ and $p^{\prime}\coloneq[0:\cdots:0:1]\in K^{\prime}\coloneq\{[\xi_{1}:\cdots:\xi_{n}]:\xi_{1}=0\}$.
+Then for every $x\in B\setminus\{p\}$, every limit point of $t_{0}^{-n}x$ is in $K^{\prime}$; and for every $x\in B\setminus K$, one has $\lim_{n}t_{0}^{-n}x=p^{\prime}$.
+Let a finite subset $F\subset\Gamma$ be given. We claim that for every $n$ there are $s_{1},\ldots,s_{n}\in\Gamma$ such that the members of
+$\{as_{i}p:a\in F,i=1,\ldots,n\}\cup\{bs_{j}p^{\prime}:b\in F,j=1,\ldots,n\}$
+are mutually distinct and that the hyperplanes $\{s_{i}K\}$ (resp.
+$\{s_{i}K^{\prime}\}$) are in general position, i.e., for every $I\subset\{1,\ldots,n\}$ the linear dimension of $\bigcap_{i\in I}s_{i}K$ is $\max\{d-|I|,0\}$ (resp.
+likewise).
+The proof is by induction on $n$.
+Suppose that the claim holds true for n1n-1 and we are given $s_{1},\ldots,s_{n-1}\in\Gamma$ that meet the claim.
+For each $a\in\Gamma$ and $q,q^{\prime}\in B$ (which may be the same), the subset $\{s\in\Gamma:sq\neq aq^{\prime}\}$ is non-empty and Zariski-open.
+For each $a\in\Gamma\setminus\{1\}$ and $q,q^{\prime}\in B$, the subset $\{s\in\Gamma:sq\neq asq^{\prime}\}$ is also non-empty and Zariski-open.
+For example, suppose by contradiction that $sq=asq^{\prime}$ for all $s\in\Gamma$.
+Then $q=aq^{\prime}$ and $sq=asq^{\prime}=asa^{-1}q$ for all $s\in G$ by Zariski-density of $\Gamma$ in $G$, but since $\bigcup_{s}s^{-1}G_{q}s=G$, where $G_{q}$ is the stabilizers of $q$, this implies that $a$ acts as identity on $B$, meaning that $a=1$.
+Also for each proper subspaces $L$ and $L^{\prime}$ the subset $\{s\in\Gamma:L^{\prime}\not\subset sL\}$ is non-empty and Zariski-open.
+Since a finite intersection of non-empty Zariski-open subsets is non-empty, there is $s_{n}$ such that $s_{1},\ldots,s_{n}$ meet the claim.
+We continue the proof and let $s_{1},\ldots,s_{n}$ be as in the previous claim.
+We choose neighborhoods $U_{i}$ of $s_{i}p$ and $V_{i}$ of $s_{i}K$ with $U_{i}\subset V_{i}$ and likewise $U_{i}^{\prime}\subset V_{i}^{\prime}$ that satisfy that
+$\{aU_{i}:a\in F,i=1,\ldots,n\}\cup\{bU_{j}^{\prime}:b\in F,j=1,\ldots,n\}$
+are mutually disjoint and that every $I\subset\{1,\ldots,n\}$ with cardinality $d$ satisfies $\bigcap_{i\in I}V_{i}=\emptyset$ and $\bigcap_{i\in I}V_{i}^{\prime}=\emptyset$.
+Then by compactness, for each $i$ there is $n_{i}\in\mathbb{N}$ large enough so that $t_{i}\coloneq s_{i}t_{0}^{n_{i}}s_{i}^{-1}$ satisfies $t_{i}^{-1}(B\setminus U_{i})\subset V_{i}^{\prime}$ and $t_{i}^{-1}(B\setminus V_{i})\subset U_{i}^{\prime}$.
+We further choose $x_{0}\in B$ and set $C_{i}\coloneq\{g\in\Gamma:gx_{0}\in U_{i}\}$ and $D_{i}\coloneq\{g\in\Gamma:gx_{0}\in V_{i}\}$.
+Thus $t_{1},\ldots,t_{n}$ and $C_{i}\subset D_{i}$ verify $\mathrm{P}_{\mathrm{PHP}}$ with
+$\sup_{x\in\Gamma}|\{i:x\in D_{i}\}\cup\{j:x\in t_{j}^{-1}(\Gamma\setminus C_{j})\}|\leq 2(d-1).$

### v7

@@
 (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$.
-Note that an extreme boundary $X$ with $|X|>2$ accommodates an axial sequence if and only if it is topologically-free.
+Note that a second countable extreme boundary $X$ with $|X|>2$ accommodates an axial sequence if and only if it is topologically-free.
 Theorem 1.
@@
 He thanks G.
-Patchell for the stimulating talk at MFO and a useful comment on this paper and E.
+Patchell for the stimulating talk at MFO and a useful comment on this paper, E.
 Breuillard and I.
-Vigdorovich for insights into $\mathrm{PSL}(d,\mathbb{R})$.
+Vigdorovich for insights into $\mathrm{PSL}(d,\mathbb{R})$, and L.
+Robert for a comment.
 2. Trees and covariant representations
@@
 Definition.
-We say that a group $\Gamma$ has property $\mathrm{P}_{\mathrm{PHP}}$ if for every finite subset $F\subset\Gamma$, $\varepsilon>0$, and $n\in\mathbb{N}$, there are $t_{i}\in\Gamma$ and $C_{i}\subset D_{i}\subset\Gamma$, $i=1,\ldots,n$, satisfying that the members of
+We say that a group $\Gamma$ has property $\mathrm{P}_{\mathrm{PHP}}$ if for every finite subset $F\subset\Gamma$ and $\varepsilon>0$, there are $n\in\mathbb{N}$ and $t_{i}\in\Gamma$ and $C_{i}\subset D_{i}\subset\Gamma$ for $1\leq i\leq n$ satisfying that the members of
 $\{aC_{i}:a\in F,\,i=1,\ldots,n\}\cup\{bt_{j}^{-1}(\Gamma\setminus D_{j}):b\in F,\,j=1,\ldots,n\}$
@@
 Proof.
-Given a finite subset $F\subset\Gamma$, $\varepsilon>0$, and $n\in\mathbb{N}$, we take $t_{i}$ and $C_{i}\subset D_{i}\subset\Gamma$ as in Definition.
+Given a finite subset $F\subset\Gamma$ and $\varepsilon>0$, we take $t_{i}$ and $C_{i}\subset D_{i}\subset\Gamma$ for $1\leq i\leq n$ as in Definition.
 Let $P_{i}\leq Q_{i}$ denote the diagonal projections in $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to the subsets $C_{i}\subset D_{i}$ and set $P_{i}^{\prime}\coloneq t_{i}^{-1}Q_{i}^{\perp}t_{i}$ and $Q_{i}^{\prime}\coloneq t_{i}^{-1}P_{i}^{\perp}t_{i}$.
@@
 Moreover, $T^{*}aT=0=\tau(a)$ for every $a\in F\setminus\{1\}$. One has by Schur’s test
-$\|\sum_{i}(Q_{i}-P_{i})t_{i}\|$ $=\|\left[\begin{matrix}Q_{1}-P_{1}&\cdots&Q_{n}-P_{n}\end{matrix}\right]\left[\begin{matrix}t_{1}&&\\ &\ddots&\\ &&t_{n}\end{matrix}\right]\left[\begin{matrix}Q^{\prime}_{1}\\ \vdots\\ Q^{\prime}_{n}\end{matrix}\right]\|$
-$\leq\|\sum_{i}Q_{i}-P_{i}\|^{1/2}\|\sum_{i}Q^{\prime}_{i}\|^{1/2}\leq\varepsilon n^{1/2}$
+$\|\sum_{i}(Q_{i}-P_{i})t_{i}\|$ $=\|\sum_{i}(Q_{i}-P_{i})t_{i}(Q_{i}^{\prime}-P_{i}^{\prime})\|\leq\|\sum_{i}Q_{i}\|^{1/2}\|\sum_{i}Q^{\prime}_{i}\|^{1/2}\leq\varepsilon n^{1/2}$
 and hence
@@
 We sketch the proof by adopting that of Lemma 4 in [dlH].
-Let a finite subset $F\subset\Gamma$ and $n$ be given.
+Let a finite subset $F\subset\Gamma$ and $\varepsilon>0$ be given.
+Take $n\geq\varepsilon^{-2}$.
 Then, there are $t_{1},\ldots,t_{n}$ in $\Gamma$ and mutually disjoint open subsets $U_{i}^{\pm}$ such that $t_{i}(L\setminus U_{i}^{-})\subset U_{i}^{+}$ for every $i$.
@@
 We fix a point $x_{0}\in L$ and set $C_{i}\coloneq D_{i}\coloneq\{s\in\Gamma:sx_{0}\in U_{i}^{+}\}$.
-Then they witness $\mathrm{P}_{\mathrm{PHP}}$ with $\varepsilon=0$.
+They witness $\mathrm{P}_{\mathrm{PHP}}$.
 ∎
@@
 Then for every $x\in B\setminus\{p\}$, every limit point of $t_{0}^{-n}x$ is in $K^{\prime}$; and for every $x\in B\setminus K$, one has $\lim_{n}t_{0}^{-n}x=p^{\prime}$.
-Let a finite subset $F\subset\Gamma$ be given. We claim that for every $n$ there are $s_{1},\ldots,s_{n}\in\Gamma$ such that the members of
+Let a finite subset $F\subset\Gamma$ and $\varepsilon>0$ be given. We claim that for every $n$ there are $s_{1},\ldots,s_{n}\in\Gamma$ such that the members of
 $\{as_{i}p:a\in F,i=1,\ldots,n\}\cup\{bs_{j}p^{\prime}:b\in F,j=1,\ldots,n\}$
@@
 We further choose $x_{0}\in B$ and set $C_{i}\coloneq\{g\in\Gamma:gx_{0}\in U_{i}\}$ and $D_{i}\coloneq\{g\in\Gamma:gx_{0}\in V_{i}\}$.
-Thus $t_{1},\ldots,t_{n}$ and $C_{i}\subset D_{i}$ verify $\mathrm{P}_{\mathrm{PHP}}$ with
-$\sup_{x\in\Gamma}|\{i:x\in D_{i}\}\cup\{j:x\in t_{j}^{-1}(\Gamma\setminus C_{j})\}|\leq 2(d-1).$
+Thus, for $n$ large enough, $t_{1},\ldots,t_{n}$ and $C_{i}\subset D_{i}$ verify $\mathrm{P}_{\mathrm{PHP}}$ with
+$\sup_{x\in\Gamma}|\{i:x\in D_{i}\}\cup\{j:x\in t_{j}^{-1}(\Gamma\setminus C_{j})\}|\leq 2(d-1)\leq\varepsilon n^{1/2}.$